### v4

# An entanglement witness and an effective criterion for bipartite states of different dimensions

Abstract
We present a simple approach for generation of a diverse set of positive maps between spaces of different dimensions. The proposed method enables the construction of Entanglement Witnesses tailored for systems in $d_{1}\times d_{2}$ dimensions. We then show how application of this entanglement witness to a general density matrix leads to a simple calculable criterion in closed form. As a support for the power of this criterion, we apply it to a variety of (one-parameter) PPT entangled states and show that the ranges of the parameters where these states are entangled are much larger than previously reported.
Introduction
Entanglement stands out as one of the most remarkable distinctions between the quantum and classical realms, playing a pivotal role in numerous quantum protocols, for review papers see [1, 2, 3]. As a result, the creation, manipulation, application, and identification of entanglement have profound significance. The identification of entangled states, in particular, presents a formidable challenge, underscored by its classification as an NP-Hard problem in mathematical terms [4]. Hence, the pursuit of viable techniques for detecting entangled states takes on paramount importance.
Entangled states are referred to as states that are not separable, and separable states refer to states that can be written as follows[5]:
$\rho=\sum p_{i}\rho_{i}^{(1)}\otimes\rho_{i}^{(2)},$ (1)
where $0\leq p_{i}\leq 1$ and $\sum p_{i}=1$. One strategy for identifying entanglement involves investigating whether a state can be expressed in the aforementioned form. This, however, poses a formidable challenge. As a result, there is a pressing need for more effective methods for the detection of entanglement. Over time, diverse techniques have emerged for this purpose. The Peres criterion is considered one of the primary methods in this regard[6]: It’s relatively straightforward to see that the partial transpose of all separable states is positive. This is seen simply by noting that for any state of the form (1), one has
$(I\otimes T)(\rho)\geq 0,$ (2)
where $T$ is the transpose, hence the name partial transpose. Therefore, if the partial transpose of a quantum state is not positive, it must be entangled. However, this criterion alone cannot identify all types of quantum entangled states because there are entangled states known as PPT, whose partial transpose are positive. In fact separable states are a subset of PPT states and only when $d_{1}d_{2}\leq 6$, these two sets are equal. Here $d_{1}$ and $d_{2}$ are the dimensions of the two quantum systems. The effectiveness of the Peres criterion (2) hinges on a basic property of the transpose map $T$ which is a positive and yet not completely positive map. Therefore it was natural to search for more general types of positive maps lacking the property of complete positivity and thus constructing more general Entanglement Witnesses than the simple partial transpose. In fact, it has been known that any entanglement witness $W$ can be written in the form
$W(M)=(I\otimes M)|\phi^{+}\rangle\langle\phi^{+}|,$
where $M$ is a positive map and $|\phi^{+}\rangle=\frac{1}{\sqrt{d}}\sum_{\mu=0}^{d-1}|\mu\mu\rangle$, is a maximally entangled state[7]. In other words, $W(M)$ is the Choi-matrix of the map $M$, which due to its lack of complete positivity can have negative eigenvalue. Such a witness has the property
$\operatorname{Tr}(W\sigma)\geq 0,\quad\forall\sigma\in S,$
where $S$ is the set of all separable states. Alternatively, one can think of $W$ as a block-positive operator, i.e.
$\langle\psi,\phi|W|\psi,\phi\rangle,\ \forall\ \psi\in H_{1}\ ,\forall\ \phi\in H_{2},$
which has negative eigenvalue when acting on the whole Hilbert space $H_{1}\otimes H_{2}$. Thus the problem of constructing entanglement witnesses becomes equivalent to constructing positive but not completely positive maps. Based on this equivalence, entanglement witnesses were introduced in many works, see [7, 8, 9, 10] and references therein. For more recent works, see [11, 12, 13]. In practical situations, where the two particles are far apart in remote laboratories, application of any entanglement witnesses found so far, requires a common reference frame between the two parties holding the particles. It was recently shown in [14] that by a process called incomplete teleportation, all these entangled states are still able to detect entanglement of remote parties, even in the absence of a common frame of reference.
In contrast with completely positive maps which due to the Kraus theorem can be easily constructed, construction of positive maps is not a straightforward task, and appropriate methods need to be proposed for this purpose. In addition to the finding new classes of witnesses, the relation of EW’s and positive maps, have led to many new insights on both [8]. For example, a basic distinction refers to decomposable witnesses which are of the form
$W=P+Q^{\Gamma},$
where $P,Q$ are positive operators and $\Gamma$ is the operator of partial transposition. Any witness that cannot be written in the above form is called indecomposable. It can be proved that decomposable witnesses cannot identify PPT entangled states. In this sense these are somewhat weaker than indecomposable witnesses. One can also characterize witnesses by their optimality and extremality [8], a subject which we will not deal with in this work. The line that we follow is inspired by a number of recent works [15, 16, 17, 18, 19, 20] in which specific types of positive maps have been defined and the corresponding witnesses have been constructed.
It is worth noting that in almost all the mentioned articles, the entanglement witnesses constructed have been designed for two systems of equal dimensions. In this paper, we propose a simple method for identification of entanglement in systems with dimensions of $d_{1}\times d_{2}$. This dimension-specific applicability enhances the versatility of the method, making it a valuable tool in different practical scenarios. Often, due to practical constraints in laboratory setups, certain measurements may be more accessible or accurate than others. Our method elegantly accommodates this by allowing for the construction of entanglement witnesses based solely on those restricted measurements which are feasible in the given experimental context(See examples).
The structure of this paper is as follows: In Section 1, we examine the main idea of this article, which is to construct a new class of positive maps. In Section 2, an alternative method is utilized to obtain the same results that presented in Section 1. In section 3 we provide a simple criterion built with our Entanglement Witness which easily detects entangled states. In Section 4, we present several examples which shows the power of our criterion in detecting PPT entangled states. We end up the paper with a conclusion.
1 Construction of positive maps for arbitrary different dimensions
Consider the Hilbert space $H_{d_{1}}$ of dimension $d_{1}$ and let a pure state $X\in D(H_{d_{1}})$, (the space of states on $H_{d_{1}}$) be written as
$X=\frac{1}{d_{1}}I_{d_{1}}+\sum_{i=1}^{d_{1}^{2}-1}x_{i}\Gamma_{i},$
where $\{\Gamma_{i},\ \ \ i=1\cdots d_{1}^{2}-1\}$ are a set of orthonormal traceless Hermitian operators $\in L(H_{d_{1}})$, i.e.
$\operatorname{Tr}(\Gamma_{i}\Gamma_{j})=\delta_{ij},\ \ \ \operatorname{Tr}\Gamma_{i}=0.$
Purity of the state $X$ (i.e. $\operatorname{Tr}X^{2}=1$) and the above normalization conditions restricts the norm of the vector ${\bf x}=(x_{1},\cdots x_{d_{1}^{2}-1})$ to
${\bf x}\cdot{\bf x}=\frac{d_{1}-1}{d_{1}},$ (3)
where ${\bf x}\cdot{\bf x}:=\sum_{i=1}^{d_{1}^{2}-1}x_{i}^{2}$.
Let $\Phi:D(H_{d_{1}})\longrightarrow D(H_{d_{2}})$ be unital and trace-preserving map which acts as follows:
$\Phi(\frac{1}{d_{1}}I_{d_{1}})=\frac{1}{d_{2}}I_{d_{2}}\ \ \ \ ,\ \ \ \Phi(\Gamma_{i})=\frac{1}{d_{2}}\sum_{k=1}^{d_{2}^{2}-1}\Lambda_{ik}\Omega_{k}.$ (4)
In fact, $\Lambda_{ik}$ are real numbers representing the characterization of the map $\Phi$. Here, our definition remains consistent with the trace-preserving property of the map, in contrast to the definition presented in [18], which is grounded in the property of such a map that takes an identity matrix in $D(H_{d_{1}})$ to a identity matrix in $D(H_{d_{2}})$. The factor $\frac{1}{d_{2}}$ in the second equation is written for later convenience, and $\{\Omega_{k},\ \ \ k=1\cdots d_{2}^{2}-1\}$ are a set of orthonormal traceless Hermitian operators $\in L(H_{d_{2}})$, i.e.
$\operatorname{Tr}(\Omega_{k}\Omega_{l})=\delta_{kl},\ \ \ \operatorname{Tr}\Omega_{k}=0.$
Then $\Phi$ maps $X$ to
$\Phi(X)=\frac{1}{d_{2}}I_{d_{2}}+\frac{1}{d_{2}}\sum_{i=1}^{d_{1}^{2}-1}\sum_{k=1}^{d_{2}^{2}-1}x_{i}\Lambda_{ik}\Omega_{k}.$ (5)
Remark 1
Up to now we have assumed that $X$ is a pure state. This can be extended to arbitrary operatores $X\in L(H_{d_{1}})$ as follows:
$\Phi(X)=\frac{\operatorname{Tr}(X)}{d_{2}}I_{d_{2}}+\frac{1}{d_{2}}\sum_{i=1}^{d_{1}^{2}-1}\sum_{k=1}^{d_{2}^{2}-1}x_{i}\Lambda_{ik}\Omega_{k}.$
To ensure positivity of the map $\Phi$, we now use Mehta’s Lemma [21] which we state below:
Mehta’s Lamma (1989): Let $A$ be a Hermitian matrix of dimension $D$, if
$\operatorname{Tr}(A^{2})\leq\frac{(\operatorname{Tr}A)^{2}}{D-1},$ (6)
then $A$ is positive.
In view of linearity of the map $\Phi$, to prove positivity of the map $\Phi$, it is sufficient to prove that $\Phi(X)$ is positive for any pure state $X$. That is the map $\Phi$ is positive if the following condition holds
$\operatorname{Tr}(\Phi(X)^{2})\leq\frac{(\operatorname{Tr}\Phi(X))^{2}}{d_{2}-1}=\frac{1}{d_{2}-1}.$ (7)
To this end, we calculate $\operatorname{Tr}(\Phi(X)^{2})$, which in view of orthonormality condition (4), turns out to be
$\operatorname{Tr}(\Phi(X)^{2})=\frac{1}{d_{2}}+\frac{1}{d_{2}^{2}}{\bf x}^{T}{\Lambda}\Lambda^{T}{\bf x},$
where the matrix $\Lambda$ is composed of $\Lambda_{ij}$ elements. In view of (6), we demand that
$\frac{1}{d_{2}}+\frac{1}{d_{2}^{2}}{\bf x}^{T}{\Lambda}\Lambda^{T}{\bf x}\leq\frac{1}{d_{2}-1},\hskip 28.45274pt\forall{\bf x}.$
We can now replace this with an even stronger inequality, namely
$\frac{1}{d_{2}}+\frac{1}{d_{2}^{2}}\lambda_{max}(\Lambda{\Lambda^{T}}){\bf x}\cdot{\bf x}\leq\frac{1}{d_{2}-1},$ (8)
where $\lambda_{max}:=\lambda_{max}(\Lambda\Lambda^{T})$, is the largest eigenvalue of the matrix $\Lambda\Lambda^{T}$, or equivalently $\lambda_{max}$ is the square of the largest singular value of the matrix $\Lambda^{T}$. Inserting (8) in (7) and using (3), yields the following constraint on $\lambda_{max}$
$\lambda_{max}\leq\frac{d_{1}d_{2}}{(d_{1}-1)(d_{2}-1)}.$ (9)
The EW corresponding to the map (5) is given by [7]
$W=(I_{d_{1}}\otimes\Phi)|\beta\rangle\langle\beta|,$
where $|\beta\rangle=\frac{1}{\sqrt{d}}\sum_{\mu=0}^{d_{1}-1}|\mu,\mu\rangle$ is a maximally entangled state. To do this calculation, we extend the orthonormal sets $\{\Gamma_{i}\}$ and $\{\Omega_{k}\}$ respectively to orthonormal bases for the space of matrices $M_{d_{1}}$ and $M_{d_{2}}$, namely to $\{\Gamma_{\mu}\}=\{\Gamma_{0}=\frac{I}{\sqrt{d_{1}}},\Gamma_{i}\}$ and $\{\Omega_{\mu}\}=\{\Omega_{0}=\frac{I}{\sqrt{d_{2}}},\Omega_{i}\}$ and taking remark 1 into account, we write
$W$ $=$ $\frac{1}{d_{1}}\sum_{\mu,\nu}|\mu\rangle\langle\nu|\otimes\Phi(|\mu\rangle\langle\nu|)$ (10)
$=$ $\frac{1}{d_{1}}\sum_{\mu,\nu}|\mu\rangle\langle\nu|\otimes\sum_{\alpha=0}^{d_{1}^{2}-1}\sum_{\beta=0}^{d_{2}^{2}-1}\Lambda_{\beta\alpha}\operatorname{Tr}(|\mu\rangle\langle\nu|\Gamma_{\beta})\Omega_{\alpha}$
$=$ $\frac{1}{d_{1}}\sum_{\alpha,\beta}\Lambda_{\beta\alpha}\Gamma^{T}_{\beta}\otimes\Omega_{\alpha}.$ (11)
where $\Lambda_{00}=\frac{1}{d_{2}},\Lambda_{0i}=0,\Lambda_{i0}=0$. After multiplication by an overall factor $d_{1}d_{2}$, this leads to the following witness
$W=I_{d_{1}}\otimes I_{d_{2}}+\sum_{i=1}^{d_{1}^{2}-1}\sum_{k=1}^{d_{2}^{2}-1}{\Lambda}_{ik}\Gamma^{T}_{i}\otimes\Omega_{k}.$
Since the set $\Gamma^{T}_{i}$ is also an orthonormal set, we can rename it to $\Gamma_{i}$. Thus, the following matrix can be a entanglement witness too:
$W=I_{d_{1}}\otimes I_{d_{2}}+\sum_{i=1}^{d_{1}^{2}-1}\sum_{k=1}^{d_{2}^{2}-1}{\Lambda}_{ik}\Gamma_{i}\otimes\Omega_{k}.$ (12)
Therefore for any two dimensions, any operator of the above form can be an entanglement witness, provided that the condition (9) holds. That condition in fact proves Block positivity of this Witness. Whether or not a $W$ constructed in this way, has also a negative eigenvalue depends on the matrix $\Lambda$ that we choose. In all previous papers that were mentioned in the introduction, $\Lambda$ should be orthogonal. These comprise a much smaller class compared to the $\Lambda$ matrices in this paper. Moreover, finding orthogonal matrices in high dimensions is not a straightforward task. In this paper, one can use any matrix with any properties to construct entanglement witnesses, as long as the largest eigenvalue of $\Lambda\Lambda^{T}$ satisfies eq. (9).
In the next section, we derive the same result of eq. (12) in an alternative way by focusing on the EW itself, rather than the positive map which leads to it via its Choi-matrix. In view of this alternative construction, one can first choose the local operators (according to the experimental restrictions) and then tuning the coefficients for defining the final entanglement witness. This is a significant advantage compared to methods where a positive map may lead to a witness which requires non-feasible measurements.
2 An alternative construction of Entanglement Witnesses
In this section we directly apply Mehta’s Lemma to a form of EW which we take as follows:
$W=I_{d_{1}}\otimes I_{d_{2}}+\sum_{i=1}^{d_{1}^{2}-1}\sum_{j=1}^{d_{2}^{2}-1}\Lambda_{ij}\Gamma_{i}\otimes\Omega_{j}.$
We demand that $W$ be block-positive. This means that for any vector $|v\rangle\in H_{d_{2}}$, the following matrix should be positive
$B\equiv\langle v|W|v\rangle=I_{d_{1}}+\sum_{i=1}^{d_{1}^{2}-1}\sum_{j=1}^{d_{2}^{2}-1}\Lambda_{ij}\Gamma_{i}\langle v|\Omega_{j}|v\rangle.$
Now we use the following:
$\operatorname{Tr}(B)=d_{1},$
and
$\operatorname{Tr}(B^{2})=d_{1}+\sum_{i,j=1}^{d_{2}^{2}-1}(\Lambda^{T}\Lambda)_{ij}\langle v|\Omega_{i}|v\rangle\langle v|\Omega_{j}|v\rangle,$
where we have used the relations $\operatorname{Tr}(\Gamma_{i})=0,\ \ \ \operatorname{Tr}(\Gamma_{i}\Gamma_{j})=\delta_{ij}$. We now resort to the Mehta’s Lemma and demand that the following condition hold for all the vectors $|v\rangle$
$\operatorname{Tr}(B^{2})\leq\frac{\operatorname{Tr}(B)^{2}}{d_{1}-1}.$
This leads to
[…]
Defining the vector ${\bf q}$ with components $q_{i}:=\langle v|\Omega_{i}|v\rangle$, this can be written as
[…]
or in a stronger form
$d_{1}+\lambda_{max}(\Lambda^{T}\Lambda){\bf q}.{\bf q}\leq\frac{d_{1}^{2}}{d_{1}-1}.$ (13)
It remains to calculate ${\bf q}.{\bf q}$, and this is easily found by noting that
${\bf q}.{\bf q}=\sum_{i=1}^{d_{2}^{2}-1}\langle v|\Omega_{i}|v\rangle\langle v|\Omega_{i}|v\rangle=\sum_{i=1}^{d_{2}^{2}-1}\langle v,v|\Omega_{i}\otimes\Omega_{i}|v,v\rangle.$
Using the identity
$\frac{1}{d}I+\sum_{i}\Omega_{i}\otimes\Omega_{i}=P,$ (14)
where $P$ is the permutation operator, we find that
${\bf q}.{\bf q}=\langle v,v|P-\frac{1}{d_{2}}I|v,v\rangle=1-\frac{1}{d_{2}}.$
Inserting this in (13), leads to the final form for the upper bound of $\lambda_{max}$, namely
$\lambda_{max}(\Lambda^{T}\Lambda)\leq\frac{d_{1}d_{2}}{(d_{1}-1)(d_{2}-1)},$ (15)
which is identical to the bound obtained by the positive map in section 1. However we should also ensure that $W$ is not trivially positive, that is we should remove the possibility that all its eigenvalues be positive. We again use Mehta’s lemma. To this end we use
$\operatorname{Tr}(W)=d_{1}d_{2},$
and
$\operatorname{Tr}(W^{2})=d_{1}d_{2}+\operatorname{Tr}(\Lambda^{T}\Lambda),$
and ensure that the following inequality be satisfied
$\operatorname{Tr}(W^{2})>\frac{\operatorname{Tr}(W)^{2}}{d_{1}d_{2}-1},$
or
$\frac{d_{1}d_{2}}{d_{1}d_{2}-1}<\operatorname{Tr}(\Lambda^{T}\Lambda).$ (16)
Combining (15) and (16), we arrive at the final condition for the parameters of the matrix $\Lambda$, in order that $W$ can be an entanglement witness:
$\frac{d_{1}d_{2}}{d_{1}d_{2}-1}<\operatorname{Tr}(\Lambda^{T}\Lambda),\ \ \ \ {\rm and}\ \ \ \lambda_{max}\leq\frac{d_{1}d_{2}}{(d_{1}-1)(d_{2}-1)}.$ (17)
Note that the second condition is sufficient for the block-positivity of $W$, while the first condition is necessary for $W$, ensuring that it is not a positive matrix.
3 A simple criterion
Consider now the form of the witness in equation (12) subject to the condition
$\lambda_{max}\leq\frac{d_{1}d_{2}}{(d_{1}-1)(d_{2}-1)},$ (18)
which is sufficient for its block-positivity. Let an arbitrary bi-partite state be expressed as follows in an orthogonal trace-less basis:
$\rho=\frac{1}{d_{1}d_{2}}I_{d_{1}}\otimes I_{d_{2}}+\sum_{i=1}^{d_{1}^{2}-1}{a}_{i}\Gamma_{i}\otimes I_{d_{2}}+\sum_{j=1}^{d_{2}^{2}-1}{b}_{j}I_{d_{1}}\otimes\Omega_{j}+\sum_{i=1}^{d_{1}^{2}-1}\sum_{j=1}^{d_{2}^{2}-1}{r}_{ij}\Gamma_{i}\otimes\Omega_{j}.$
It is readily seen that
$\operatorname{Tr}(W\rho)=1+\sum_{ij}\Lambda_{ij}r_{ij}=1+\operatorname{Tr}(R\Lambda^{T}),$ (19)
where the matrix $R$ is composed of $r_{ij}$ elements. In order to make $\operatorname{Tr}(W\rho)<0$, which at the same time proves that $W$ is indeed an entanglement witness (i.e. has a negative eigenvalue) and detects the entanglement of $\rho$, we try to choose $\Lambda$ in such a way to minimize $tr(R\Lambda^{T})$, while maintaining the condition (18). To achieve this, we first express the matrix $R$ using singular value decomposition:
$R=USV.$
Here we have used the fact that $R$ is a real matrix, which implies the matrices $U\in M_{d_{1}\times d_{1}}$ and $V\in M_{d_{2}\times d_{2}}$ are real orthogonal matrices. The matrix $S\in M_{d_{1}\times d_{2}}$ is a rectangular diagonal matrix with non-negative entries on the main diagonal. Now we choose the matrix $\Lambda^{T}$ as follows:
$\Lambda^{T}=V^{T}LU^{T},$
where $L\in M_{d_{2}\times d_{1}}$ is chosen to be a rectangular diagonal matrix whose entries on its main diagonal are equal to $l_{i}=-\sqrt{\frac{d_{1}d_{2}}{(d_{1}-1)(d_{2}-1)}}$. This leads to
$\Lambda^{T}\Lambda=V^{T}(LL^{T})V,$
where $LL^{T}\in M_{d_{2}\times d_{2}}$ is a matrix whose diagonal elements are either zero or equal to $\frac{d_{1}d_{2}}{(d_{1}-1)(d_{2}-1)}$. This means that the condition (9) on the maximum eigenvalue of $\Lambda^{T}\Lambda$ is satisfied. We then find that
$\operatorname{Tr}(\Lambda^{T}R)=\operatorname{Tr}(SL)=-\sqrt{\frac{d_{1}d_{2}}{(d_{1}-1)(d_{2}-1)}}\operatorname{Tr}(S).$
In this way we have obtained the lowest value of $\operatorname{Tr}(\Lambda^{T}R)$, while maintaining the condition (18) of block-positivity of $W$. Therefore, if this value is less than $-1$, which in view of (19) is equivalent to $\operatorname{Tr}(W\rho)<0$, we conclude that the state $\rho$ is entangled. In addition, one can show that $Tr(S)=Tr(\sqrt{RR^{T}})$. Therefore, we have the following simple criterion:
${\rm If\ \ }\gamma:=\operatorname{Tr}(S)>\sqrt{\frac{(d_{1}-1)(d_{2}-1)}{d_{1}d_{2}}},\ \ \ {\rm the\ state\ is\ entangled.}$ (20)
The condition that we derived does not require finding the elements of the matrix $\Lambda$, and only requires recasting the density matrix in terms of a basis chosen by orthogonal trace-less matrices, and calculating the singular values of the matrix $R$.
It should be mentioned that the singular values of $R$ and hence the value of $\gamma$ do not depend on the bases which we choose for the two parts for expanding the density matrix $\rho$, as long as these bases are orthonormal and traceless.
Remark: Interestingly, this criterion is a special case of Theorem 1 in [22], where the authors show that this is a powerful criterion and unifies several criteria known before.
We will now show by several examples that this criterion, originated from the EW defined in (12), is indeed powerful enough to detect entanglement of a class of PPT entangled states which has not been previously detected.
4 Examples
In this section, we consider a few examples and show the power of our criterion. In all the following examples, except example 4.2, the states are PPT entangled states reported in previous works. We will show that the criterion given in (20), can detect entanglement in a larger interval of the parameters previously reported, and it also can detect PPT entangled states with different dimensions, i.e. $d_{1}\neq d_{2}$. For checking the following examples, the reader should note that we have used the following order for Gellmann matrices:
[…]
$\textrm{for }\ \ \ \ \ \ \frac{1}{2}d(d-1)\leq i<d(d-1):\ \ \ \ \ \ \Gamma_{i}\in\Gamma_{y},$ (21)
$\textrm{for }\ \ \ \ \ \ d(d-1)\leq i<d^{2}:\hskip 28.45274pt\hskip 28.45274pt\ \ \Gamma_{i}\in\Gamma_{z},$ (22)
where for $1\leq m<n\leq d$,
$\Gamma_{x}=\big{\{}\ \ \frac{1}{\sqrt{2}}(|m\rangle\langle n|+|n\rangle\langle m|)\ \big{\}},$
$\Gamma_{y}=\big{\{}\ \ \frac{i}{\sqrt{2}}(|m\rangle\langle n|-|n\rangle\langle m|)\ \ \big{\}},$
and for $11\leq n\leq d-$ 1
$\Gamma_{z}=\big{\{}\frac{1}{\ \ \sqrt{n(n+1)}}(\sum_{m=0}^{n-1}|m\rangle\langle m|-n|n\rangle\langle n|)\ \ \big{\}}.$
Moreover in all those cases where the state is a PPT state and its partial transpose equals to itself, i.e. $\rho=\rho^{\Gamma}$, we can use the anti-symmetricity of the Gell-Mann matrices in $\Gamma_{y}$ to simplify the calculations in the following way:
Let at least one of the matrices $\Gamma_{i}$ or $\Omega_{j}$ (say $\Gamma_{i}$) be anti-symmetric, then we have
$r_{ij}=\operatorname{Tr}(\rho(\Gamma_{i}\otimes\Omega_{j}))=\operatorname{Tr}(\rho^{\Gamma}(\Gamma_{i}\otimes\Omega_{j}))=\operatorname{Tr}(\rho.(\Gamma_{i}^{T}\otimes\Omega_{j}))=-\operatorname{Tr}(\rho(\Gamma_{i}\otimes\Omega_{j}))=0.$
The same is true when $\Omega_{j}$ is anti-symmetric.
4.1 Example 1:
In this example, we demonstrate that the method explained in this paper can identify PPT entangled states, and thus, it can generate indecomposable EWs. It has been shown in [23] that the following state is PPT entangled state when $x\neq 1$ (for $x=1$ this is a separable state):
$\rho=\frac{1}{N}\left(\begin{array}[]{ccc|ccc|ccc}1&.&.&.&1&.&.&.&1\\ .&\frac{1}{x}&.&.&.&.&.&.&.\\ .&.&x&.&.&.&.&.&.\\ \hline\cr.&.&.&x&.&.&.&.&.\\ 1&.&.&.&1&.&.&.&1\\ .&.&.&.&.&\frac{1}{x}&.&.&.\\ \hline\cr.&.&.&.&.&.&\frac{1}{x}&.&.\\ .&.&.&.&.&.&.&x&.\\ 1&.&.&.&1&.&.&.&1\\ \end{array}\right).$ (23)
Here dots represent zeros, $N=3+x+x^{-1}$ is the normalization factor and $x>0$. The above state can be written in the basis of normalized Gell-Mann matrices as follows:
[…]
$+$ $\left(1-\frac{\xi}{2}\right)(\Gamma_{7}\otimes\Gamma_{7}+\Gamma_{8}\otimes\Gamma_{8})-\frac{\sqrt{3}}{2}\eta(\Gamma_{7}\otimes\Gamma_{8}-\Gamma_{8}\otimes\Gamma_{7})\Big{]},$
where $\xi=x+x^{-1}$ and $\eta=x-x^{-1}$. The eigenvalues of the matrix $R$ are easily found to be
$\lambda(R)=\frac{1}{N}\{1(3),-1(3),(1-\frac{\xi}{2})\pm i\frac{\sqrt{3}}{2}\eta\},$
where the numbers in parentheses indicate the multiplicity of each singular value. Thus, taking into account the normalization factor $N$, we find
$Tr(S)=\frac{6+2\sqrt{(1-\frac{\xi}{2})^{2}+\frac{3}{4}\eta^{2}}}{3(1+\xi)}.$
Condition (20) for entanglement of the state now amounts to
$\frac{1}{4}(\xi^{2}-\eta^{2})<\xi-1,$
or after simplification (inserting $\xi=x+x^{-1}$ and $\eta=x-x^{{}_{1}}$) leads to the following condition for the state (23) to be entangled:
$2<x+x^{-1}.$
This shows that the state $\rho_{x}$ is entangled for all values of $x$, except for $x=1$, which coincides with the result of [23].
4.2 Example 2:
This example demonstrates that our method can identify the entanglement of all NPT states in the following form in any dimension $d$:
$\rho=p|\psi\rangle\langle\psi|+(1-p)\frac{I}{d},\quad|\psi\rangle=\frac{1}{\sqrt{d}}\sum_{i=0}^{d-1}|i,i\rangle.$
To calculate $\gamma$ in equation (20), we need to obtain the coefficients $r_{ij}$:
$r_{ij}=p\operatorname{Tr}[\frac{1}{d}(\sum_{k,l}|kk\rangle\langle ll|).(\Gamma_{i}\otimes\Gamma_{j})]=\frac{p}{d}\operatorname{Tr}(\Gamma_{i}\Gamma_{j}^{T})=\pm\frac{p}{d}\delta_{ij}.$
Here we have noticed that the expression $\operatorname{Tr}(\Gamma_{i}\Gamma_{j}^{T})$ equals $\delta_{ij}$ when $\Gamma_{j}^{T}=\Gamma_{j}^{\dagger}$ and equals $-\delta_{ij}$ when $\Gamma_{j}^{\dagger}=-\Gamma_{j}^{T}$. As it is evident, this state is diagonal in the Gell-Mann matrix basis, and the absolute value of all its singular values is equal to $\frac{p}{d}$, and the number of its singular values is also equal to $d^{2}-1$. Therefore, we find for this state
$\gamma=\operatorname{Tr}(S)=\frac{p}{d}(d^{2}-1),$
which, according to eq (20), should be greater than $\frac{d-1}{d}$ so the state is entangled. Thus the range of $p$ for the state to be entangled is $p>\frac{1}{d+1}$. This is the same value obtained from the Peres criterion which in this case follows from
$\rho^{\Gamma}=(1-p)\frac{I}{d^{2}}+p\frac{1}{d}F,$
where $F$ is the swap operator. The eigenstates are the symmetric and antisymmetric states and the smallest eigenvalue, namely $\frac{1-p}{d^{2}}-\frac{p}{d}$, is negative when $p>\frac{1}{d+1}$ and hence the state is entangled. For this example the two methods give equivalent results. This result has already been reported in [24].
4.3 Example 3
In this example, we provide a PPT entangled state which can be detected by the criterion in equation (20). To construct a PPT entangled state we define the following Unextendible Product Basis in $C_{3}\otimes C_{4}:$
$|\phi_{1}\rangle=\frac{1}{\sqrt{2}}|0\rangle|1-2\rangle,\quad|\phi_{2}\rangle=\frac{1}{\sqrt{6}}|0\rangle(2|0\rangle-|1+2\rangle),\quad|\phi_{3}\rangle=\frac{1}{\sqrt{2}}|1\rangle|1-2\rangle,\quad$
$|\phi_{4}\rangle=\frac{1}{\sqrt{2}}|2\rangle|1-2\rangle,\quad|\phi_{5}\rangle=\frac{1}{\sqrt{6}}|2\rangle(|1+2\rangle-2|3\rangle),\quad|\phi_{6}\rangle=\frac{1}{\sqrt{2}}|1-2\rangle|0\rangle,\quad$ (25)
$|\phi_{7}\rangle=\frac{1}{\sqrt{2}}|0-1\rangle|3\rangle,\quad|\phi_{8}\rangle=\frac{1}{\sqrt{12}}|0+1+2\rangle|0+1+2+3\rangle,$ (26)
and construct the PPT entangled state $\rho=I-\sum_{i}|\phi_{i}\rangle\langle\phi_{i}|$ [8], whose explicit form is given by
$\rho=\frac{1}{48}\left(\begin{array}[]{c|c|c|c}3\ \epsilon&-\epsilon&-\epsilon&-\epsilon\\ \hline\cr-\epsilon&D&D^{\prime}&-\epsilon\\ \hline\cr-\epsilon&D^{\prime T}&D&-\epsilon\\ \hline\cr-\epsilon&-\epsilon&-\epsilon&3\ \epsilon\\ \end{array}\right),$
where $\epsilon$ is the 3 dimensional square matrix all of its elements are equal to 1, and
$D=\left(\begin{array}[]{ccc}5&-1&-1\\ -1&5&-1\\ -1&-1&5\end{array}\right),\quad D^{\prime}=\left(\begin{array}[]{ccc}-1&5&-1\\ -1&-1&5\\ 5&-1&-1\end{array}\right).$
The matrix $R$ is simply obtained by noting that its elements are given by $r_{mn}=\operatorname{Tr}(\rho(\Gamma_{m}\otimes\Omega_{n}))=-\sum_{i}\langle\phi_{i}|\Gamma_{m}\otimes\Omega_{n}|\phi_{i}\rangle$. The result is
$R=\frac{1}{24}\left(\begin{array}[]{c|c|c}-\epsilon_{3\times 6}&0_{3\times 6}&A_{3\times 3}\\ \hline\cr 0_{3\times 6}&0_{3\times 6}&0_{3\times 3}\\ \hline\cr B_{2\times 6}&0_{2\times 6}&C_{2\times 3}\end{array}\right),$
where $\epsilon_{3\times 6}$ is a $3\times 6$ matrix all of whose elements are 1, and similarly ${0}_{3\times 6}$ is a matrix all of whose elements are zero. Also
$A$ $=$ $\left(\begin{array}[]{ccc}0&0&-3\sqrt{\frac{3}{2}}\\ 0&0&0\\ 3&\sqrt{3}&\sqrt{\frac{3}{2}}\end{array}\right),$ (30)
$B$ $=$ $\left(\begin{array}[]{cccccc}{2}&{2}&-1&0&0&0\\ \frac{2}{\sqrt{3}}&\frac{2}{\sqrt{3}}&\frac{1}{\sqrt{3}}&0&\frac{-4}{\sqrt{3}}&\frac{-4}{\sqrt{3}}\end{array}\right),$ (33)
$C$ $=$ $\left(\begin{array}[]{ccc}0&0&-\frac{1}{2}\sqrt{\frac{3}{2}}\\ -\frac{1}{\sqrt{3}}&-\frac{1}{3}&-\frac{5}{6\sqrt{2}}\end{array}\right).$ (36)
The singular values of this matrix are obtained numerically and one finds that
$\gamma=\operatorname{Tr}(S)=1.429>\sqrt{\frac{3\times 4}{2\times 3}}=\sqrt{2},$
which shows that while $\rho$ is a PPT entangled state, it can be detected by the criterion (20).
4.4 Example 4
In [23], among other things, the following family of PPT entangled states in $C_{4}\times C_{4}$ are introduced
$\rho=\frac{1}{N}\left(\begin{array}[]{cccc|cccc|cccc|cccc}1&.&.&.&.&1&.&.&.&.&1&.&.&.&.&1\\ .&\frac{1}{a}&.&.&.&.&.&.&.&.&.&.&.&.&1&.\\ .&.&1&.&.&.&.&.&1&.&.&.&.&.&.&.\\ .&.&.&a&.&.&1&.&.&.&.&.&.&.&.&.\\ \hline\cr.&.&.&.&a&.&.&.&.&.&.&1&.&.&.&.\\ 1&.&.&.&.&1&.&.&.&.&1&.&.&.&.&1\\ .&.&.&1&.&.&\frac{1}{a}&.&.&.&.&.&.&.&.&.\\ .&.&.&.&.&.&.&1&.&.&.&.&.&1&.&.\\ \hline\cr.&.&1&.&.&.&.&.&1&.&.&.&.&.&.&.\\ .&.&.&.&.&.&.&.&.&a&.&.&1&.&.&.\\ 1&.&.&.&.&1&.&.&.&.&1&.&.&.&.&1\\ .&.&.&.&1&.&.&.&.&.&.&\frac{1}{a}&.&.&.&.\\ \hline\cr.&.&.&.&.&.&.&.&.&1&.&.&\frac{1}{a}&.&.&.\\ .&.&.&.&.&.&.&1&.&.&.&.&.&1&.&.\\ .&1&.&.&.&.&.&.&.&.&.&.&.&.&a&.\\ 1&.&.&.&.&1&.&.&.&.&1&.&.&.&.&1\\ \end{array}\right),$ (37)
where $N=4a+\frac{4}{a}+8$, and it is shown that their entanglement can be detected by an entanglement witness, called $W_{ext}$, which shows that these states are entangled in the interval $10<a<$ 1. We now show that the interval of entanglement of these states is in fact much larger and covers the whole positive line $0<a<\infty$, except the point $a=1$. To show this, the $R$ matrix corresponding to (37) is calculated, which turns out to be
$R=\frac{1}{(a+1)^{2}}\big{(}A\oplus A^{\prime}\oplus B\big{)},$
where
$A=\left(\begin{array}[]{cccccc}\frac{a}{4}&.&.&.&.&\frac{a}{4}\\ .&\frac{a}{2}&.&.&.&.\\ .&.&\frac{a}{4}&\frac{a}{4}&.&.\\ .&.&\frac{a}{4}&\frac{a}{4}&.&.\\ .&.&.&.&\frac{a}{2}&.\\ \frac{a}{4}&.&.&.&.&\frac{a}{4}\end{array}\right),\ \ \ \ \ A^{\prime}=\left(\begin{array}[]{cccccc}-\frac{a}{4}&.&.&.&.&\frac{a}{4}\\ .&.&.&.&.&.\\ .&.&-\frac{a}{4}&-\frac{a}{4}&.&.\\ .&.&-\frac{a}{4}&-\frac{a}{4}&.&.\\ .&.&.&.&.&.\\ \frac{a}{4}&.&.&.&.&-\frac{a}{4}\end{array}\right),$
and
$B=\left(\begin{array}[]{ccc}-\frac{(a-1)^{2}}{8}&-\frac{a^{2}+2a-3}{8\sqrt{3}}&-\frac{(a-1)a}{2\sqrt{6}}\\ \frac{3a^{2}-2a-1}{8\sqrt{3}}&-\frac{(a-1)^{2}}{24}&-\frac{a^{2}+a-2}{6\sqrt{2}}\\ \frac{a-1}{2\sqrt{6}}&\frac{2a^{2}-a-1}{6\sqrt{2}}&-\frac{(a-1)^{2}}{12}\end{array}\right),$
with singular values given as
$\Big{\{}\frac{\left|a-1\right|}{4\left|a+1\right|}\ (2),\ \frac{\left|a\right|}{2(a+1)^{2}}\ (6),\ \frac{(a-1)^{2}}{4(a+1)^{2}}\ (1)\Big{\}}.$
This leads to
$\gamma=\frac{1}{4}\frac{(a-1)^{2}}{(a+1)^{2}}+3\frac{|a|}{(a+1)^{2}}+\frac{1}{2}\left|\frac{a-1}{a+1}\right|.$
Figure 1 shows that the value of $\gamma$ is always larger than $\sqrt{\frac{(d_{1}-1)(d_{2}-1)}{d_{1}d_{2}}}=\frac{3}{4}$ except for $a=1$.
Figure 1: Using the criterion intruduced in this paper, i.e. eq (17), we can see that state (37) is entangled for $0<a<\infty;a\neq 1$, which is much larger the range intruduced in [23].
Note that by $a<\infty$, we are emphasizing that $a$ should be a finite value. In fact, in the limit $a\longrightarrow\infty$, the state approaches to
$\rho_{\infty}=\frac{1}{4}\Big{(}|03\rangle\langle 03|+|10\rangle\langle 10|+|11\rangle\langle 11|+|22\rangle\langle 22|\Big{)},$
which is obviously separable. Incidentally, this is the limiting case where $\gamma$ approaches the value $\frac{3}{4}$.
4.5 Example 5
This example is a continuation of the previous example and demonstrates that for the second PPT entangled state introduced in [23], the interval of entanglement (indicated as $10<b<$ 1), is indeed much larger and as in the previous example covers all positive and finite values of $b$ different from $1$. The state is given by
$\rho=\frac{1}{N}\left(\begin{array}[]{cccc|cccc|cccc|cccc}1&.&.&.&.&1&.&.&.&.&1&.&.&.&.&1\\ .&\frac{1}{b}&.&.&1&.&.&.&.&.&.&.&.&.&.&.\\ .&.&1&.&.&.&.&.&1&.&.&.&.&.&.&.\\ .&.&.&b&.&.&.&.&.&.&.&.&1&.&.&.\\ \hline\cr.&1&.&.&b&.&.&.&.&.&.&.&.&.&.&.\\ 1&.&.&.&.&1&.&.&.&.&1&.&.&.&.&1\\ .&.&.&.&.&.&\frac{1}{b}&.&.&1&.&.&.&.&.&.\\ .&.&.&.&.&.&.&1&.&.&.&.&.&1&.&.\\ \hline\cr.&.&1&.&.&.&.&.&1&.&.&.&.&.&.&.\\ .&.&.&.&.&.&1&.&.&b&.&.&.&.&.&.\\ 1&.&.&.&.&1&.&.&.&.&1&.&.&.&.&1\\ .&.&.&.&.&.&.&.&.&.&.&\frac{1}{b}&.&.&1&.\\ \hline\cr.&.&.&1&.&.&.&.&.&.&.&.&\frac{1}{b}&.&.&.\\ .&.&.&.&.&.&.&1&.&.&.&.&.&1&.&.\\ .&.&.&.&.&.&.&.&.&.&.&1&.&.&b&.\\ 1&.&.&.&.&1&.&.&.&.&1&.&.&.&.&1\\ \end{array}\right),$
where $N=4b+\frac{4}{b}+8$, and its corresponding matrix is
$R=\frac{1}{(1+b)^{2}}\left(\frac{b}{2}I_{6}\oplus 0_{6}\oplus A\right),$
where $I_{6}$ and $0_{6}$ are respectively the identity matrix and the zero matrix in $C_{6}$ and
$A=\left(\begin{array}[]{ccc}-\frac{1}{8}(b-1)^{2}&-\frac{b^{2}+2b-3}{8\sqrt{3}}&-\frac{(b-1)b}{2\sqrt{6}}\\ \frac{3b^{2}-2b-1}{8\sqrt{3}}&-\frac{1}{24}(b-1)^{2}&-\frac{b^{2}+b-2}{6\sqrt{2}}\\ \frac{b-1}{2\sqrt{6}}&\frac{2b^{2}-b-1}{6\sqrt{2}}&-\frac{1}{12}(b-1)^{2}\\ \end{array}\right),$
with singular values given as
$\Big{\{}\frac{\left|b-1\right|}{4\left|b+1\right|}\ (2),\ \frac{\left|b\right|}{2(b+1)^{2}}\ (6),\ \frac{(b-1)^{2}}{4(b+1)^{2}}\ (1)\Big{\}},$
which is the same as the previous example and gives the same result.
4.6 Example 6
In [25], Terhal introduced a family of UPB for $3\times n$ systems. Here we take $n=4$ and show that our criterion can detect the PPT entangled state built by them in these dimensions.
The UPB states for the $3\times 4$ system are:
$|\psi_{0}\rangle$ $=$ $\frac{1}{\sqrt{2}}|0\rangle\otimes(|1\rangle-|3\rangle),$
$|\psi_{1}\rangle$ $=$ $\frac{1}{\sqrt{2}}|1\rangle\otimes(|2\rangle-|3\rangle),$ (38)
$|\psi_{2}\rangle$ $=$ $\frac{1}{\sqrt{2}}|2\rangle\otimes(|0\rangle-|3\rangle),$ (39)
$|\psi_{3}\rangle$ $=$ $\frac{1}{\sqrt{2}}(|0\rangle-|1\rangle)\otimes|0\rangle,$ (40)
$|\psi_{4}\rangle$ $=$ $\frac{1}{\sqrt{2}}(|1\rangle-|2\rangle)\otimes|1\rangle,$ (41)
$|\psi_{5}\rangle$ $=$ $\frac{1}{\sqrt{2}}(|2\rangle-|0\rangle)\otimes|2\rangle,$ (42)
$|\psi_{6}\rangle$ $=$ $\frac{1}{\sqrt{12}}\sum_{i=0}^{2}\sum_{j=0}^{3}|i\rangle\otimes|j\rangle.$ (43)
The PPT entangled state $\rho=I-\sum_{0}^{6}|\psi_{i}\rangle\langle\psi_{i}|$ turns out to be
$\rho=\frac{1}{60}\left(\begin{array}[]{ccc|ccc|ccc|ccc}5&-1&-1&-1&5&-1&-1&-1&-1&-1&-1&-1\\ -1&5&-1&5&-1&-1&-1&-1&-1&-1&-1&-1\\ -1&-1&5&-1&-1&-1&-1&-1&-1&-1&5&-1\\ \hline\cr-1&5&-1&5&-1&-1&-1&-1&-1&-1&-1&-1\\ 5&-1&-1&-1&5&-1&-1&-1&-1&-1&-1&-1\\ -1&-1&-1&-1&-1&5&-1&-1&-1&5&-1&-1\\ \hline\cr-1&-1&-1&-1&-1&-1&5&5&-1&-1&-1&-1\\ -1&-1&-1&-1&-1&-1&5&5&-1&-1&-1&-1\\ -1&-1&-1&-1&-1&-1&-1&-1&5&-1&-1&5\\ \hline\cr-1&-1&-1&-1&-1&5&-1&-1&-1&5&-1&-1\\ -1&-1&5&-1&-1&-1&-1&-1&-1&-1&5&-1\\ -1&-1&-1&-1&-1&-1&-1&-1&5&-1&-1&5\\ \end{array}\right).$ (56)
The corresponding $R$ matrix is given by
$R=\frac{1}{30}\left(\begin{array}[]{c|c|c}-\epsilon_{3\times 6}&0_{3\times 6}&A\\ \hline\cr 0_{3\times 6}&0_{3\times 6}&0_{3\times 3}\\ \hline\cr B&0_{2\times 6}&0_{2\times 3}\end{array}\right),$
where
$A=\left(\begin{array}[]{ccc}3&\sqrt{3}&\sqrt{\frac{3}{2}}\\ 0&-2\sqrt{3}&\sqrt{\frac{3}{2}}\\ -3&-\sqrt{3}&\sqrt{\frac{3}{2}}\end{array}\right),$
and
$B=\left(\begin{array}[]{cccccc}0&0&0&0&3&-3\\ 0&0&0&-2\sqrt{3}&\sqrt{3}&\sqrt{3}\\ \end{array}\right).$
The singular values of $R$ are found to be
$\left\{\frac{1}{2\sqrt{10}},\frac{1}{5\sqrt{2}},\frac{1}{5\sqrt{2}},\frac{1}{5\sqrt{2}},\frac{1}{5\sqrt{2}}\right\},$
and we find
$\gamma=\frac{1}{2\sqrt{10}}+\frac{4}{5\sqrt{2}}=\big{(}\frac{8+\sqrt{5}}{10}\big{)}\frac{1}{\sqrt{2}},$
which is greater than $\sqrt{\frac{(d_{1}-1)(d_{2}-1)}{d_{1}d_{2}}}=\frac{1}{\sqrt{2}}$. This shows that the PPT state (56) is entangled.
5 Conclusion
In this paper, a class of Entanglement Witnesses for two particle systems with different dimensions was presented. Specifically we showed in section (3), how the use of this class of EW’s can be reduced to a simple criterion, (equation (20)), which can be checked on the form of the density matrix. We then showed by checking several one-parameter family of PPT entangled states that this criterion can detect their entanglement in a region of parameters much larger than previously thought.

### v5

@@
-An entanglement witness and an effective criterion for bipartite states of different dimensions
+Positive maps and Entanglement Witnesses in different dimensions
 Abstract
-We present a simple approach for generation of a diverse set of positive maps between spaces of different dimensions.
-The proposed method enables the construction of Entanglement Witnesses tailored for systems in $d_{1}\times d_{2}$ dimensions.
-We then show how application of this entanglement witness to a general density matrix leads to a simple calculable criterion in closed form.
-As a support for the power of this criterion, we apply it to a variety of (one-parameter) PPT entangled states and show that the ranges of the parameters where these states are entangled are much larger than previously reported.
+We present a continuous, multiparameter family of positive maps between spaces of differing dimensions.
+This framework facilitates the construction of Entanglement Witnesses (EWs) specifically designed for systems in $d_{1}\times d_{2}$ dimensions.
+We derive a simple, closed-form criterion for detecting entanglement in general density matrices based on these witnesses.
+To demonstrate the effectiveness of this criterion, we apply it to a range of Positive Partial Transpose (PPT) entangled states, revealing that the parameter regions where these states exhibit entanglement are larger than previously reported.
+Furthermore, we prove that non-unital EWs, corresponding to non-unital maps, are not more powerful than unital EWs, thus supporting the focus on unital positive maps in recent studies.
+Our method complements existing approaches to separability criteria for density matrices in different dimensions.
 Introduction
-Entanglement stands out as one of the most remarkable distinctions between the quantum and classical realms, playing a pivotal role in numerous quantum protocols, for review papers see [1, 2, 3].
-As a result, the creation, manipulation, application, and identification of entanglement have profound significance.
+Any conceivable operation on a quantum system, be it a measurement or noisy evolution, can be represented by a Completely Positive Trace-preserving (CPT) map or a quantum channel for short.
+A map of this kind not only maps density matrices to density matrices, but all of its extensions also have this property.
+The name of completely positive map comes from this last property.
+At first sight it seems, due to this stringent completely positive property, that construction of CPT maps is much more difficult than that of simply positive maps.
+However it has turned out that it is quite the contrary.
+CPT maps can readily be constructed, among other methods, by resorting to their Kraus representation [1].
+In fact any such map from a system $A$ to another system $B$ (possibly with different dimension can be written as
+${\cal E}(\rho)=\sum_{\alpha=1}^{K}A_{\alpha}\rho A_{\alpha}^{\dagger}$ (1)
+where $A_{\alpha}:H_{A}\longrightarrow H_{B}$ are arbitrary linear operators whose only condition is the trace-preserving property, $\sum_{\alpha=1}A_{\alpha}^{\dagger}A_{\alpha}=I_{A}$, a condition which can be easily fullfiled, i.e.
+one simply chooses arbitrary operators $A_{\alpha},\ \ \alpha=1,\cdots K-1$ and by multiplying them with suitable scale factors ensures that $\sum_{\alpha=1}^{K-1}A_{\alpha}^{\dagger}A_{\alpha}\leq I_{A}$ and defines the last Kraus operator simply as $A_{K}:=\sqrt{I-\sum_{\alpha=1}^{K-1}A_{\alpha}^{\dagger}A_{\alpha}}$.
+While construction of instances of CPT maps is not challenging, classification of these maps, specially in higher dimensions, and deciding whether or not a given map is CPT is generally difficult.
+On the other hand construction of instances of maps which are only positive, but not completely positive, is not straightforward at all and as expected, their classification has remained an open problem despite decades of efforts by physicists and mathematicians.
+The need for construction of instances of positive maps arises from their role in entanglement theory [2, 3, 4].
 The identification of entangled states, in particular, presents a formidable challenge, underscored by its classification as an NP-Hard problem in mathematical terms [5].
@@
 where $0\leq p_{i}\leq 1$ and $\sum p_{i}=1$.
-One strategy for identifying entanglement involves investigating whether a state can be expressed in the aforementioned form.
-This, however, poses a formidable challenge.
+It is not easy at all to determine whether or not a given state can be expressed in the aforementioned form.
 As a result, there is a pressing need for more effective methods for the detection of entanglement.
 Over time, diverse techniques have emerged for this purpose.
-The Peres criterion is considered one of the primary methods in this regard[6]: It’s relatively straightforward to see that the partial transpose of all separable states is positive.
-This is seen simply by noting that for any state of the form (1), one has
+The Peres criterion is considered one of the primary methods in this regard[7], i.e.
+a state which is separable satisifes
 $(I\otimes T)(\rho)\geq 0,$ (3)
@@
 Here $d_{1}$ and $d_{2}$ are the dimensions of the two quantum systems.
-The effectiveness of the Peres criterion (2) hinges on a basic property of the transpose map $T$ which is a positive and yet not completely positive map.
-Therefore it was natural to search for more general types of positive maps lacking the property of complete positivity and thus constructing more general Entanglement Witnesses than the simple partial transpose.
-In fact, it has been known that any entanglement witness $W$ can be written in the form
-$W(M)=(I\otimes M)|\phi^{+}\rangle\langle\phi^{+}|,$
-where $M$ is a positive map and $|\phi^{+}\rangle=\frac{1}{\sqrt{d}}\sum_{\mu=0}^{d-1}|\mu\mu\rangle$, is a maximally entangled state[7].
-In other words, $W(M)$ is the Choi-matrix of the map $M$, which due to its lack of complete positivity can have negative eigenvalue.
-Such a witness has the property
+In higher dimensions one should search for other criteria of entanglement, many of which have been reviewed in [8].
+The most important ones are called Entanglement Witnesses (EW), see [9] for a review.
+An EW refers to a Hermitian operator $W$, which has negative eigenvalues but has the following property:
 $\operatorname{Tr}(W\sigma)\geq 0,\quad\forall\sigma\in S,$
-where $S$ is the set of all separable states. Alternatively, one can think of $W$ as a block-positive operator, i.e.
+where $S$ is the set of all separable states.
+Therefore, a negative expectation value for $W$ on a state $\rho$, then means that $\rho$ is entangled.
+Alternatively, one can think of $W$ as a block-positive operator, i.e.
 $\langle\psi,\phi|W|\psi,\phi\rangle,\ \forall\ \psi\in H_{1}\ ,\forall\ \phi\in H_{2},$
 which has negative eigenvalue when acting on the whole Hilbert space $H_{1}\otimes H_{2}$.
+Here comes the important link between entanlement witnesses and positive maps.
+In fact, it has been known that any entanglement witness $W$ can be written in the form
+$W=(I\otimes\Phi)|\phi^{+}\rangle\langle\phi^{+}|,$
+where $\Phi$ is a positive (but not completely positive) maps, and $|\phi^{+}\rangle=\frac{1}{\sqrt{d}}\sum_{\mu=0}^{d-1}|\mu\mu\rangle$, is a maximally entangled state[10].
+In other words, $W$ is the Choi-matrix of the map $\Phi$, which due to its lack of complete positivity can have negative eigenvalue.
 Thus the problem of constructing entanglement witnesses becomes equivalent to constructing positive but not completely positive maps.
+This is an important and known result which connects the long-standing problem of classification of positive maps in pure mathematics to the new problem of detection of entanglement quantum information theory.
 Based on this equivalence, entanglement witnesses were introduced in many works, see [10, 9, 11, 8] and references therein.
-For more recent works, see [11, 12, 13].
-In practical situations, where the two particles are far apart in remote laboratories, application of any entanglement witnesses found so far, requires a common reference frame between the two parties holding the particles.
-It was recently shown in [14] that by a process called incomplete teleportation, all these entangled states are still able to detect entanglement of remote parties, even in the absence of a common frame of reference.
-In contrast with completely positive maps which due to the Kraus theorem can be easily constructed, construction of positive maps is not a straightforward task, and appropriate methods need to be proposed for this purpose.
-In addition to the finding new classes of witnesses, the relation of EW’s and positive maps, have led to many new insights on both [8].
-For example, a basic distinction refers to decomposable witnesses which are of the form
-$W=P+Q^{\Gamma},$
-where $P,Q$ are positive operators and $\Gamma$ is the operator of partial transposition.
-Any witness that cannot be written in the above form is called indecomposable.
-It can be proved that decomposable witnesses cannot identify PPT entangled states.
-In this sense these are somewhat weaker than indecomposable witnesses.
-One can also characterize witnesses by their optimality and extremality [8], a subject which we will not deal with in this work.
-The line that we follow is inspired by a number of recent works [15, 16, 17, 18, 19, 20] in which specific types of positive maps have been defined and the corresponding witnesses have been constructed.
-It is worth noting that in almost all the mentioned articles, the entanglement witnesses constructed have been designed for two systems of equal dimensions.
-In this paper, we propose a simple method for identification of entanglement in systems with dimensions of $d_{1}\times d_{2}$.
-This dimension-specific applicability enhances the versatility of the method, making it a valuable tool in different practical scenarios.
-Often, due to practical constraints in laboratory setups, certain measurements may be more accessible or accurate than others.
-Our method elegantly accommodates this by allowing for the construction of entanglement witnesses based solely on those restricted measurements which are feasible in the given experimental context(See examples).
-The structure of this paper is as follows: In Section 1, we examine the main idea of this article, which is to construct a new class of positive maps.
-In Section 2, an alternative method is utilized to obtain the same results that presented in Section 1.
-In section 3 we provide a simple criterion built with our Entanglement Witness which easily detects entangled states.
+For more recent works in this direction, see [12, 13, 14], where the authors generally start from a restricted class of positive maps, mostly between spaces of equal dimensions and construct the corresponding EW’s.
+These maps are always uniital and their corresponding affine transformation is a rotation performed by an orthogonal matrix.
+In this work, we construct a large class of positive maps which firstly act between spaces of different dimensions $d_{1}\times d_{2}$.
+Besides detecting entanglement in heterogenuous systems, this makes it possible to witness entanglement in a multi-partite system, for example when $k$ qubits are entnagled to $n-k$ qubits in a many-body system.
+In particular, we
+a) provide sufficient conditions for our maps to be positive, but not completely positive, so that their Choi matrix is an EW, and then
+b) provide sufficient conditions for our maps to be completely positive. Both criteria are expressed in terms of the affine transformation corresponding to this map.
+c) We then prove that among the EW’s constructed from positive maps, the unital EW’s, that is, those constructed from unital maps are stronger than those corresponding to non-unital maps.
+Thus we rigorously support the focus on unital positive maps in recent studies [12, 13, 14].
+d) We reduce the previous considerations to a simple criterion which solely depends on the density matrix itself and does not require any type of entanglement witnesses.
+Thus simply by checking this critria, one can decide whether or not a given state is entangled.
+e) By studying several examples, we show that the EW’s generated in this way, can detect entanglement of states for a larger range of parameters than previously found.
+f) Finally in an appendix, we restrict ourselves to unital EW’s and show how by using the concept of block-positivity, we can directly construct the EW’s derived from the positive maps.
+The structure of this paper is as follows: In Section 1, we set up our notation.
+Then in section 2, we state Mehta’s lemma and its proof which is the basis of all our construction.
+In section 3, we use this lemma [15] to construct general positive maps between spaces of different dimensions.
+By general maps, we mean maps which are not necessarily unital.
+This part extends the previous works which have constructed EW’s along similar lines [16, 17, 18].
+In section 4 we reduce the previous considerations to a simple criterion which solely depends on the density matrix itself and does not require any type of entanglement witnesses.
+We also show in this section that in each family of maps, the unital ones lead to the strongest type of witnesses.
 In Section 5, we present several examples which shows the power of our criterion in detecting PPT entangled states.
 We end up the paper with a conclusion.
-1 Construction of positive maps for arbitrary different dimensions
-Consider the Hilbert space $H_{d_{1}}$ of dimension $d_{1}$ and let a pure state $X\in D(H_{d_{1}})$, (the space of states on $H_{d_{1}}$) be written as
-$X=\frac{1}{d_{1}}I_{d_{1}}+\sum_{i=1}^{d_{1}^{2}-1}x_{i}\Gamma_{i},$
-where $\{\Gamma_{i},\ \ \ i=1\cdots d_{1}^{2}-1\}$ are a set of orthonormal traceless Hermitian operators $\in L(H_{d_{1}})$, i.e.
-$\operatorname{Tr}(\Gamma_{i}\Gamma_{j})=\delta_{ij},\ \ \ \operatorname{Tr}\Gamma_{i}=0.$
-Purity of the state $X$ (i.e. $\operatorname{Tr}X^{2}=1$) and the above normalization conditions restricts the norm of the vector ${\bf x}=(x_{1},\cdots x_{d_{1}^{2}-1})$ to
-${\bf x}\cdot{\bf x}=\frac{d_{1}-1}{d_{1}},$ (3)
-where ${\bf x}\cdot{\bf x}:=\sum_{i=1}^{d_{1}^{2}-1}x_{i}^{2}$.
-Let $\Phi:D(H_{d_{1}})\longrightarrow D(H_{d_{2}})$ be unital and trace-preserving map which acts as follows:
-$\Phi(\frac{1}{d_{1}}I_{d_{1}})=\frac{1}{d_{2}}I_{d_{2}}\ \ \ \ ,\ \ \ \Phi(\Gamma_{i})=\frac{1}{d_{2}}\sum_{k=1}^{d_{2}^{2}-1}\Lambda_{ik}\Omega_{k}.$ (4)
-In fact, $\Lambda_{ik}$ are real numbers representing the characterization of the map $\Phi$.
-Here, our definition remains consistent with the trace-preserving property of the map, in contrast to the definition presented in [18], which is grounded in the property of such a map that takes an identity matrix in $D(H_{d_{1}})$ to a identity matrix in $D(H_{d_{2}})$.
-The factor $\frac{1}{d_{2}}$ in the second equation is written for later convenience, and $\{\Omega_{k},\ \ \ k=1\cdots d_{2}^{2}-1\}$ are a set of orthonormal traceless Hermitian operators $\in L(H_{d_{2}})$, i.e.
-$\operatorname{Tr}(\Omega_{k}\Omega_{l})=\delta_{kl},\ \ \ \operatorname{Tr}\Omega_{k}=0.$
-Then $\Phi$ maps $X$ to
-$\Phi(X)=\frac{1}{d_{2}}I_{d_{2}}+\frac{1}{d_{2}}\sum_{i=1}^{d_{1}^{2}-1}\sum_{k=1}^{d_{2}^{2}-1}x_{i}\Lambda_{ik}\Omega_{k}.$ (5)
-Remark 1
-Up to now we have assumed that $X$ is a pure state. This can be extended to arbitrary operatores $X\in L(H_{d_{1}})$ as follows:
-$\Phi(X)=\frac{\operatorname{Tr}(X)}{d_{2}}I_{d_{2}}+\frac{1}{d_{2}}\sum_{i=1}^{d_{1}^{2}-1}\sum_{k=1}^{d_{2}^{2}-1}x_{i}\Lambda_{ik}\Omega_{k}.$
-To ensure positivity of the map $\Phi$, we now use Mehta’s Lemma [21] which we state below:
-Mehta’s Lamma (1989): Let $A$ be a Hermitian matrix of dimension $D$, if
+In an appendix, section 8, an alternative method based on the concept of block-positiviy is used to obtain the same results derived in section 3.
+1 Notations and Conventions
+Let $H_{d}$ be the Hilbert space of dimension $d$ and $L(H_{d})$ the space of linear operators acting on $H_{d}$.
+The half space of positive operators is denoted by $L(H_{d})^{+}$.
+We will be dealing with maps $\Phi:L^{+}(H_{d_{1}})\longrightarrow L^{+}(H_{d_{2}})$.
+The Hilbert-Schmidt norm of a real matrix $A$ is denoted by $\|A\|=\operatorname{Tr}(A^{T}A)^{1/2}$.
+By $S\in M_{d_{1}\times d_{2}}$, we mean a rectangular matrix of dimension $d_{1}\times d_{2}$, by $M\in O(d)$, we mean an orthogonal square matrix of dimension $d$.
+Let $A$ be a rectangular matrix.
+Then by $a_{max}$ we mean the largest eigenvalue of the positive symmetric matrix $\sqrt{A^{T}A}$ or equivalently $\sqrt{AA^{T}}$.
+A basis of orthonormal Hermitian and traceless matrices acting on $H_{d_{1}}$ is denoted by $\{\Gamma_{i},\ i=1\cdots d_{1}^{2}-1\}$ and the same basis of matrices acting on $H_{d_{2}}$ is denoted by $\{\Omega_{\mu},\mu=1,\cdots d_{2}^{2}-1\}$.
+Therefore we have $\operatorname{Tr}(\Gamma_{i})=\operatorname{Tr}(\Omega_{\mu})=0$ and $\operatorname{Tr}(\Gamma_{i}\Gamma_{j})=\delta_{ij}$ and $\operatorname{Tr}(\Omega_{\mu}\Omega_{\nu})=\delta_{\mu,\nu}.$ Finally we use the convention of summing over repeated indices.
+In all formulas, $i$ runs from $1$ to $d_{1}^{2}-1$ and $\mu$ from $1$ to $d_{2}^{2}-1$.
+The state of a quantum system is denoted by $\rho\in D(H_{d})$, where $D(H_{d})\subset L(H_{d})^{+}$ is the convex subset of positive operators with unit trace.
+Let $X$ be a positive operator in $L^{+}(H_{d_{1}})$ which can be expanded as
+$X={x_{0}}I_{d_{1}}+x_{i}\Gamma_{i}=x_{0}I_{d_{1}}+{\bf x}\cdot\bm{\Gamma},$ (4)
+where $(x_{0},{\bf x}=(x_{1},\cdots x_{d_{1}^{2}-1}))$ are real parameters.
+For any positive operator $X$, it necessarily holds that $\operatorname{Tr}(X^{2})\leq\operatorname{Tr}(X)^{2}$ which leads to
+${\bf x}\cdot{\bf x}\leq x_{0}^{2}d_{1}(d_{1}-1).$ (5)
+where ${\bf x}\cdot{\bf x}:=x_{i}x_{i}$. If the operator $X$ is pure, that is, if $X=|\psi\rangle\langle\psi|$ for a non-normalized vector $|\psi\rangle$, then the equality sign holds in (5).
+Any operator $Y$ in $L(H_{d_{2}})$ is written as
+$Y={y_{0}}I_{d_{2}}+y_{\mu}\Omega_{\mu}=y_{0}I_{d_{2}}+{\bf y}\cdot\bm{\Omega}.$ (6)
+Let $\Phi:D(H_{d_{1}})\longrightarrow D(H_{d_{2}})$ be a map which acts as follows:
+$\Phi(I_{d_{1}})=(I_{d_{2}}+t_{\mu}\Omega_{\mu})\ \ \ \ ,\ \ \ \Phi(\Gamma_{i})=\Lambda_{\mu i}\Omega_{\mu},$ (7)
+Here $t_{\mu}$ and $\Lambda_{\mu i}$ are real numbers which characterize the map $\Phi$.
+In the special case where ${\bf t}=0$, the map $\Phi$ is unital.
+Hereafter we use summation convention for summing over repeated indices.
+With these conventions, $\Phi$ maps $X$ to
+$\Phi(X)={x_{0}}I_{d_{2}}+(x_{0}t_{\mu}+x_{i}\Lambda_{\mu i})\Omega_{\mu}=x_{0}I_{d_{2}}+({\bf t}x_{0}+\Lambda{\bf x})\cdot\Omega.$ (8)
+With our convention, the map $\Phi$ corresponds to the following affine tranformation, $\Phi:(x_{0},{\bf x})\longrightarrow(x_{0},{\bf t}x_{0}+\Lambda{\bf x})$ which can be written in matrix form .
+$\Phi:\begin{pmatrix}x_{0}\\ {\bf x}\end{pmatrix}\longrightarrow\begin{pmatrix}1&{\bf 0}\\ {\bf t}&\Lambda\end{pmatrix}\begin{pmatrix}x_{0}\\ {\bf x}\end{pmatrix}.$ (9)
+where ${\bf t}$ is a $d_{2}$ dimensional vector and $\Lambda\in M_{d_{2}\times d_{1}}$.
+2 Mehta’s Lamma
+Let $A$ be a Hermitian matrix of dimension $D$, if
 $\operatorname{Tr}(A^{2})\leq\frac{(\operatorname{Tr}A)^{2}}{D-1},$ (10)
 then $A$ is positive.
-In view of linearity of the map $\Phi$, to prove positivity of the map $\Phi$, it is sufficient to prove that $\Phi(X)$ is positive for any pure state $X$.
-That is the map $\Phi$ is positive if the following condition holds
-$\operatorname{Tr}(\Phi(X)^{2})\leq\frac{(\operatorname{Tr}\Phi(X))^{2}}{d_{2}-1}=\frac{1}{d_{2}-1}.$ (7)
-To this end, we calculate $\operatorname{Tr}(\Phi(X)^{2})$, which in view of orthonormality condition (4), turns out to be
-$\operatorname{Tr}(\Phi(X)^{2})=\frac{1}{d_{2}}+\frac{1}{d_{2}^{2}}{\bf x}^{T}{\Lambda}\Lambda^{T}{\bf x},$
-where the matrix $\Lambda$ is composed of $\Lambda_{ij}$ elements. In view of (6), we demand that
-$\frac{1}{d_{2}}+\frac{1}{d_{2}^{2}}{\bf x}^{T}{\Lambda}\Lambda^{T}{\bf x}\leq\frac{1}{d_{2}-1},\hskip 28.45274pt\forall{\bf x}.$
-We can now replace this with an even stronger inequality, namely
-$\frac{1}{d_{2}}+\frac{1}{d_{2}^{2}}\lambda_{max}(\Lambda{\Lambda^{T}}){\bf x}\cdot{\bf x}\leq\frac{1}{d_{2}-1},$ (8)
-where $\lambda_{max}:=\lambda_{max}(\Lambda\Lambda^{T})$, is the largest eigenvalue of the matrix $\Lambda\Lambda^{T}$, or equivalently $\lambda_{max}$ is the square of the largest singular value of the matrix $\Lambda^{T}$.
-Inserting (8) in (7) and using (3), yields the following constraint on $\lambda_{max}$
-$\lambda_{max}\leq\frac{d_{1}d_{2}}{(d_{1}-1)(d_{2}-1)}.$ (9)
+Note that Mehta’s lemma provides only a sufficient condition.
+Its converse is not true, e.g.
+there are positive matrices which do not satisfy (10), as is readily seen from an example of a diagonal matrix $A=diag(1,2,x),\ \forall\ x\geq 6$.
+Proof: The proof is taken from refernce [19].
+Let $A$ be a Hermitian matrix of size $D$ and let $\alpha={\rm Tr}A/\sqrt{{\rm Tr}A^{2}}$.
+If $\alpha\geq\sqrt{D-1}$ then $A$ is positive.
+We calculate the traces in the eigenbasis of $A$.
+Let $(x_{1},\dots x_{D})$ denote the spectrum of $A$.
+Assume first that one eigenvalue, say $x_{1}$, is negative.
+Making use of the right hand side of the standard estimation between the ${l}_{1}$ and ${l}_{2}$–norms (with prefactor $1$) of a $D$–vector, $||A||_{2}\leq||A||_{1}\leq D||A||_{2}$, we infer
+${\rm TrA}=\sum_{i=1}^{D}x_{i}<\sum_{i=2}^{D}x_{i}\leq\sqrt{D-1}\Bigl{(}\sum_{i=2}^{D}x_{i}^{2}\Bigr{)}^{1/2}<\sqrt{D-1}\sqrt{{\rm Tr}A^{2}}\ .$ (11)
+This implies that $\alpha<\sqrt{D-1}$. Hence if the opposite is true and $\alpha\geq\sqrt{D-1}$ then none of the eigenvalues $x_{i}$ could be negative, so $A\geq 0$. $\square$.
+3 Positive maps between spaces of different dimensions
+Given a quantum map $\Phi$ characterized by the affine transformation on the generalized Bloch vector ${\bf r}\longrightarrow\Lambda{\bf r}+{\bf t}$, our goal is to find the condition on this affine map which gurantees positivity of $\Phi$.
+This is stated in the following:
+3.1 Sufficient condition for positivity
+Proposition: A map $\Phi$, implementing the affine transformatoin (9) is positive if the following condition holds:
+$\|{\bf t}\|+\sqrt{d_{1}(d_{1}-1)}\lambda_{max}\leq\sqrt{\frac{d_{2}}{d_{2}-1}},$ (12)
+where $\lambda_{max}$ is the largest eigenvalue of the positive matrix $\sqrt{\Lambda\Lambda^{T}}.$
+Proof: Due to linearity of the map, it suffices to prove the statement only for pure states. For such states (see (5)), we have
+$\operatorname{Tr}(\Phi(X))=x_{0}d_{2}$
+and
+$\operatorname{Tr}(\Phi(X)^{2})=x_{0}^{2}d_{2}+\|{\bf t}x_{0}+\Lambda{\bf x}\|^{2}={x_{0}^{2}}d_{2}+x_{0}^{2}{\bf t}\cdot{\bf t}+2x_{0}{\bf t}\cdot\Lambda{\bf x}+\|\Lambda{\bf x}\|^{2}.$
+In view of the Mehta lemma, the map $\Phi$ is positive if for all $(x_{0},{\bf x})$ subject to the condition (10), the following inequality holds
+${x_{0}^{2}}d_{2}+x_{0}^{2}{\bf t}\cdot{\bf t}+2x_{0}{\bf t}\cdot\Lambda{\bf x}+\|\Lambda{\bf x}\|^{2}\leq\frac{x_{0}^{2}d_{2}^{2}}{d_{2}-1}.$ (13)
+We demand a slightly stronger condition, (at a price of loosing some of the positive maps). To this end, we note that
+$\|\Lambda{\bf x}\|^{2}\leq\lambda_{max}^{2}{\bf x}\cdot{\bf x}=\lambda_{max}^{2}x_{0}^{2}d_{1}(d_{1}-1),$ (14)
+where $\lambda_{max}^{2}$ is the largest eigenvalue of the positive matrix $\Lambda^{T}\Lambda$. We also use Cauchy-Schwartz inequality to write
+${\bf t}\cdot\Lambda{\bf x}\leq\|{\bf t}\|\ \|\Lambda{\bf x}\|\leq\|{\bf t}\|\lambda_{max}\|{\bf x}\|=\|{\bf t}\|\lambda_{max}{x_{0}}\sqrt{d_{1}(d_{1}-1)},$ (15)
+Inserting (14) and (15) in (13) and dividing both sides by $x_{0}^{2}$, we find the following sufficient condition for positivity of the map $\Phi$
+$\|{\bf t}\|^{2}+2\lambda_{max}\|{\bf t}\|\sqrt{d_{1}(d_{1}-1)}+\lambda_{max}^{2}d_{1}(d_{1}-1)\leq\frac{d_{2}}{(d_{2}-1)}.$ (16)
+Taking the square root of both sides leads to (12) which concludes the proof. $\square$
+For unital maps where ${\bf t}=0$, the above condition simplifies to
+$\lambda_{max}\leq\sqrt{\frac{d_{2}}{d_{1}(d_{1}-1)(d_{2}-1)}}.$
+3.2 Sufficient condition for complete positivity
+A map $\Phi$ is completely positive if all its extensions are also positive. We now derive a sufficient condition on the affine map for complete positivity of the corresponding map.
+Proposition: A positive map (8), implementing the affine transformatoin (9) is a completely positive map if the following condition holds:
+${\bf t}\cdot{\bf t}+d_{1}\operatorname{Tr}(\Lambda\Lambda^{T})\leq\frac{d_{2}}{d_{1}d_{2}-1}.$ (17)
+Proof: According to a celebrated theorem of Choi [20], a map is completely positive if its Choi matirx defined as below is positive
+$C_{\Phi}$ $=$ $(I_{d_{1}}\otimes\Phi)(|\phi^{+}\rangle\langle\phi^{+}|)=(|i\rangle\langle j|\otimes\Phi(|i\rangle\langle j|)$ (18)
+$=$ $|i\rangle\langle j|\otimes\Big{[}\frac{1}{d_{1}}\delta_{ij}(I_{d_{2}}+{\bf t}\cdot\Omega)+\langle j|\Gamma_{l}|i\rangle\Lambda_{\mu l}\Omega_{\mu}\Big{]}$ (19)
+$=$ $\frac{1}{d_{1}}I_{d_{1}}\otimes(I_{d_{2}}+{\bf t}\cdot\Omega)+\Lambda_{\mu l}\Gamma_{l}^{T}\otimes\Omega_{\mu}.$ (20)
+from which the traces are easily derived as
+$\operatorname{Tr}(C_{\Phi})=d_{2},\hskip 28.45274pt\operatorname{Tr}(C_{\Phi}^{2})=\frac{d_{2}+{\bf t}\cdot{\bf t}}{d_{1}}+\operatorname{Tr}(\Lambda\Lambda^{T})$ (21)
+Inserting these in Mehta’s Lemma (10), leads to (17) and completes the proof $\square$.
+Example: Maps between qubits: As an example consider the qubit maps $\Phi$ with the affine transformation (9) where $\Lambda=\begin{pmatrix}\lambda_{1}&&\\ &\lambda_{2}&\\ &&\lambda_{3}\end{pmatrix}$ and ${\bf t}=(t_{1},t_{2},t_{3})^{T}$.
+Then the condition for positivity of the map is given from (12) to be
+$\sqrt{t_{1}^{2}+t_{2}^{2}+t_{3}^{2}}+\sqrt{2}\lambda_{max}\leq\sqrt{2},$ (22)
+and the condition of complete positivity is given from (17) to be
+$\sum_{i}(t_{i}^{2}+2\lambda_{i}^{2})\leq\frac{2}{3}.$ (23)
+We shoud emphasize that since the Mehta’s condition is a sufficient conditions, the above inequalities determine only a subset of positive and completely positive maps.
+For example, if we restrict ourselves to unital maps, then the condition (17) is given by
+$\lambda_{1}^{2}+\lambda_{2}^{2}+\lambda_{3}^{2}\leq\frac{1}{3},$
+which is a sphere of radius $r=\frac{1}{\sqrt{3}}$ and is a subset of the Fujiwara-Algoet tetrahedron
+$|\lambda_{1}+\lambda_{2}|\leq 1+\lambda_{3}\ \ \ \ \ |\lambda_{1}-\lambda_{2}|\leq 1-\lambda_{3}.$
+A comparison of the volumes of these two sets shows that $30$ percents of the completely positive unital maps for qubits are detected by this simple use of Mehta’s criteria, figure (1).
+Figure 1: Example of qubit maps, with eigenvalues of affine transformation given by $\lambda_{1},\lambda_{2}$ and $\lambda_{3}$.
+Restricting to unital maps, the tethdrahedron shows the complete set of unital maps as derived from the Fujiwara-Algoet condition, the sphere shows those maps which are derived from Mehta’s condition.
+Its volume is approximately 30 percent of the volume of the tethrahedron.
+4 Entanglement Witnesses from positive maps
+It is known that there is a one-to-one correspondence between positive map which are not completely positive on the one hand and Entanglement Witnesses on the other.
 The EW corresponding to the map (8) is given by [10]
 $W=(I_{d_{1}}\otimes\Phi)|\beta\rangle\langle\beta|,$
-where $|\beta\rangle=\frac{1}{\sqrt{d}}\sum_{\mu=0}^{d_{1}-1}|\mu,\mu\rangle$ is a maximally entangled state.
-To do this calculation, we extend the orthonormal sets $\{\Gamma_{i}\}$ and $\{\Omega_{k}\}$ respectively to orthonormal bases for the space of matrices $M_{d_{1}}$ and $M_{d_{2}}$, namely to $\{\Gamma_{\mu}\}=\{\Gamma_{0}=\frac{I}{\sqrt{d_{1}}},\Gamma_{i}\}$ and $\{\Omega_{\mu}\}=\{\Omega_{0}=\frac{I}{\sqrt{d_{2}}},\Omega_{i}\}$ and taking remark 1 into account, we write
-$W$ $=$ $\frac{1}{d_{1}}\sum_{\mu,\nu}|\mu\rangle\langle\nu|\otimes\Phi(|\mu\rangle\langle\nu|)$ (10)
-$=$ $\frac{1}{d_{1}}\sum_{\mu,\nu}|\mu\rangle\langle\nu|\otimes\sum_{\alpha=0}^{d_{1}^{2}-1}\sum_{\beta=0}^{d_{2}^{2}-1}\Lambda_{\beta\alpha}\operatorname{Tr}(|\mu\rangle\langle\nu|\Gamma_{\beta})\Omega_{\alpha}$
-$=$ $\frac{1}{d_{1}}\sum_{\alpha,\beta}\Lambda_{\beta\alpha}\Gamma^{T}_{\beta}\otimes\Omega_{\alpha}.$ (11)
-where $\Lambda_{00}=\frac{1}{d_{2}},\Lambda_{0i}=0,\Lambda_{i0}=0$. After multiplication by an overall factor $d_{1}d_{2}$, this leads to the following witness
-$W=I_{d_{1}}\otimes I_{d_{2}}+\sum_{i=1}^{d_{1}^{2}-1}\sum_{k=1}^{d_{2}^{2}-1}{\Lambda}_{ik}\Gamma^{T}_{i}\otimes\Omega_{k}.$
-Since the set $\Gamma^{T}_{i}$ is also an orthonormal set, we can rename it to $\Gamma_{i}$. Thus, the following matrix can be a entanglement witness too:
-$W=I_{d_{1}}\otimes I_{d_{2}}+\sum_{i=1}^{d_{1}^{2}-1}\sum_{k=1}^{d_{2}^{2}-1}{\Lambda}_{ik}\Gamma_{i}\otimes\Omega_{k}.$ (12)
+where $|\beta\rangle=\frac{1}{\sqrt{d_{1}}}\sum_{i=0}^{d_{1}-1}|i,i\rangle$ is a maximally entangled state.
+Direct calculation, after multiplication by an overall factor $d_{1}d_{2}$, leads to the following witness
+$W=I_{d_{1}}\otimes I_{d_{2}}+t_{\mu}I_{d_{1}}\otimes\Omega_{\mu}+d_{1}{\Lambda}_{\mu i}\Gamma_{i}\otimes\Omega_{\mu}.$ (24)
+In the last step, we have used the freedom to replace the matrices $\Gamma_{i}^{T}$ by $\Gamma_{i}$, since both sets can be chosen as orthonormal set for the space of Hermitian traceless matrices on $H_{d_{1}}$.
+Given a general bi-partite density matrix as
+$\rho=\frac{1}{d_{1}d_{2}}I_{d_{1}}\otimes I_{d_{2}}+{a}_{i}\Gamma_{i}\otimes I_{d_{2}}+{b}_{\mu}I_{d_{1}}\otimes\Omega_{\mu}+{r}_{i\mu}\Gamma_{i}\otimes\Omega_{\mu},$ (25)
+one finds that
+$\operatorname{Tr}(W\rho)=1+d_{1}{\bf b}\cdot{\bf t}+d_{1}\operatorname{Tr}(R\Lambda).$ (26)
+Those EW’s which correspond to unital positive maps (i.e.
+${\bf t}=0$) are called unital entanglement witness.
 Therefore for any two dimensions, any operator of the above form can be an entanglement witness, provided that the condition (12) holds.
-That condition in fact proves Block positivity of this Witness.
-Whether or not a $W$ constructed in this way, has also a negative eigenvalue depends on the matrix $\Lambda$ that we choose.
+Whether or not an EW constructed in this way, has also a negative eigenvalue depends on the matrix $\Lambda$ that we choose.
 In all previous papers that were mentioned in the introduction, $\Lambda$ should be orthogonal.
@@
 Moreover, finding orthogonal matrices in high dimensions is not a straightforward task.
-In this paper, one can use any matrix with any properties to construct entanglement witnesses, as long as the largest eigenvalue of $\Lambda\Lambda^{T}$ satisfies eq.
-(9).
-In the next section, we derive the same result of eq.
-(12) in an alternative way by focusing on the EW itself, rather than the positive map which leads to it via its Choi-matrix.
-In view of this alternative construction, one can first choose the local operators (according to the experimental restrictions) and then tuning the coefficients for defining the final entanglement witness.
-This is a significant advantage compared to methods where a positive map may lead to a witness which requires non-feasible measurements.
-2 An alternative construction of Entanglement Witnesses
-In this section we directly apply Mehta’s Lemma to a form of EW which we take as follows:
-$W=I_{d_{1}}\otimes I_{d_{2}}+\sum_{i=1}^{d_{1}^{2}-1}\sum_{j=1}^{d_{2}^{2}-1}\Lambda_{ij}\Gamma_{i}\otimes\Omega_{j}.$
-We demand that $W$ be block-positive. This means that for any vector $|v\rangle\in H_{d_{2}}$, the following matrix should be positive
-$B\equiv\langle v|W|v\rangle=I_{d_{1}}+\sum_{i=1}^{d_{1}^{2}-1}\sum_{j=1}^{d_{2}^{2}-1}\Lambda_{ij}\Gamma_{i}\langle v|\Omega_{j}|v\rangle.$
-Now we use the following:
-$\operatorname{Tr}(B)=d_{1},$
-and
-$\operatorname{Tr}(B^{2})=d_{1}+\sum_{i,j=1}^{d_{2}^{2}-1}(\Lambda^{T}\Lambda)_{ij}\langle v|\Omega_{i}|v\rangle\langle v|\Omega_{j}|v\rangle,$
-where we have used the relations $\operatorname{Tr}(\Gamma_{i})=0,\ \ \ \operatorname{Tr}(\Gamma_{i}\Gamma_{j})=\delta_{ij}$.
-We now resort to the Mehta’s Lemma and demand that the following condition hold for all the vectors $|v\rangle$
-$\operatorname{Tr}(B^{2})\leq\frac{\operatorname{Tr}(B)^{2}}{d_{1}-1}.$
-This leads to
-[…]
-Defining the vector ${\bf q}$ with components $q_{i}:=\langle v|\Omega_{i}|v\rangle$, this can be written as
-[…]
-or in a stronger form
-$d_{1}+\lambda_{max}(\Lambda^{T}\Lambda){\bf q}.{\bf q}\leq\frac{d_{1}^{2}}{d_{1}-1}.$ (13)
-It remains to calculate ${\bf q}.{\bf q}$, and this is easily found by noting that
-${\bf q}.{\bf q}=\sum_{i=1}^{d_{2}^{2}-1}\langle v|\Omega_{i}|v\rangle\langle v|\Omega_{i}|v\rangle=\sum_{i=1}^{d_{2}^{2}-1}\langle v,v|\Omega_{i}\otimes\Omega_{i}|v,v\rangle.$
-Using the identity
-$\frac{1}{d}I+\sum_{i}\Omega_{i}\otimes\Omega_{i}=P,$ (14)
-where $P$ is the permutation operator, we find that
-${\bf q}.{\bf q}=\langle v,v|P-\frac{1}{d_{2}}I|v,v\rangle=1-\frac{1}{d_{2}}.$
-Inserting this in (13), leads to the final form for the upper bound of $\lambda_{max}$, namely
-$\lambda_{max}(\Lambda^{T}\Lambda)\leq\frac{d_{1}d_{2}}{(d_{1}-1)(d_{2}-1)},$ (15)
-which is identical to the bound obtained by the positive map in section 1.
-However we should also ensure that $W$ is not trivially positive, that is we should remove the possibility that all its eigenvalues be positive.
-We again use Mehta’s lemma.
-To this end we use
-$\operatorname{Tr}(W)=d_{1}d_{2},$
-and
-$\operatorname{Tr}(W^{2})=d_{1}d_{2}+\operatorname{Tr}(\Lambda^{T}\Lambda),$
-and ensure that the following inequality be satisfied
-$\operatorname{Tr}(W^{2})>\frac{\operatorname{Tr}(W)^{2}}{d_{1}d_{2}-1},$
-or
-$\frac{d_{1}d_{2}}{d_{1}d_{2}-1}<\operatorname{Tr}(\Lambda^{T}\Lambda).$ (16)
-Combining (15) and (16), we arrive at the final condition for the parameters of the matrix $\Lambda$, in order that $W$ can be an entanglement witness:
-$\frac{d_{1}d_{2}}{d_{1}d_{2}-1}<\operatorname{Tr}(\Lambda^{T}\Lambda),\ \ \ \ {\rm and}\ \ \ \lambda_{max}\leq\frac{d_{1}d_{2}}{(d_{1}-1)(d_{2}-1)}.$ (17)
-Note that the second condition is sufficient for the block-positivity of $W$, while the first condition is necessary for $W$, ensuring that it is not a positive matrix.
-3 A simple criterion
-Consider now the form of the witness in equation (12) subject to the condition
-$\lambda_{max}\leq\frac{d_{1}d_{2}}{(d_{1}-1)(d_{2}-1)},$ (18)
-which is sufficient for its block-positivity. Let an arbitrary bi-partite state be expressed as follows in an orthogonal trace-less basis:
-$\rho=\frac{1}{d_{1}d_{2}}I_{d_{1}}\otimes I_{d_{2}}+\sum_{i=1}^{d_{1}^{2}-1}{a}_{i}\Gamma_{i}\otimes I_{d_{2}}+\sum_{j=1}^{d_{2}^{2}-1}{b}_{j}I_{d_{1}}\otimes\Omega_{j}+\sum_{i=1}^{d_{1}^{2}-1}\sum_{j=1}^{d_{2}^{2}-1}{r}_{ij}\Gamma_{i}\otimes\Omega_{j}.$
-It is readily seen that
-$\operatorname{Tr}(W\rho)=1+\sum_{ij}\Lambda_{ij}r_{ij}=1+\operatorname{Tr}(R\Lambda^{T}),$ (19)
-where the matrix $R$ is composed of $r_{ij}$ elements.
-In order to make $\operatorname{Tr}(W\rho)<0$, which at the same time proves that $W$ is indeed an entanglement witness (i.e.
-has a negative eigenvalue) and detects the entanglement of $\rho$, we try to choose $\Lambda$ in such a way to minimize $tr(R\Lambda^{T})$, while maintaining the condition (18).
-To achieve this, we first express the matrix $R$ using singular value decomposition:
-$R=USV.$
-Here we have used the fact that $R$ is a real matrix, which implies the matrices $U\in M_{d_{1}\times d_{1}}$ and $V\in M_{d_{2}\times d_{2}}$ are real orthogonal matrices.
-The matrix $S\in M_{d_{1}\times d_{2}}$ is a rectangular diagonal matrix with non-negative entries on the main diagonal.
-Now we choose the matrix $\Lambda^{T}$ as follows:
-$\Lambda^{T}=V^{T}LU^{T},$
-where $L\in M_{d_{2}\times d_{1}}$ is chosen to be a rectangular diagonal matrix whose entries on its main diagonal are equal to $l_{i}=-\sqrt{\frac{d_{1}d_{2}}{(d_{1}-1)(d_{2}-1)}}$. This leads to
-$\Lambda^{T}\Lambda=V^{T}(LL^{T})V,$
-where $LL^{T}\in M_{d_{2}\times d_{2}}$ is a matrix whose diagonal elements are either zero or equal to $\frac{d_{1}d_{2}}{(d_{1}-1)(d_{2}-1)}$.
-This means that the condition (9) on the maximum eigenvalue of $\Lambda^{T}\Lambda$ is satisfied.
-We then find that
-$\operatorname{Tr}(\Lambda^{T}R)=\operatorname{Tr}(SL)=-\sqrt{\frac{d_{1}d_{2}}{(d_{1}-1)(d_{2}-1)}}\operatorname{Tr}(S).$
-In this way we have obtained the lowest value of $\operatorname{Tr}(\Lambda^{T}R)$, while maintaining the condition (18) of block-positivity of $W$.
-Therefore, if this value is less than $-1$, which in view of (19) is equivalent to $\operatorname{Tr}(W\rho)<0$, we conclude that the state $\rho$ is entangled.
-In addition, one can show that $Tr(S)=Tr(\sqrt{RR^{T}})$.
-Therefore, we have the following simple criterion:
-${\rm If\ \ }\gamma:=\operatorname{Tr}(S)>\sqrt{\frac{(d_{1}-1)(d_{2}-1)}{d_{1}d_{2}}},\ \ \ {\rm the\ state\ is\ entangled.}$ (20)
-The condition that we derived does not require finding the elements of the matrix $\Lambda$, and only requires recasting the density matrix in terms of a basis chosen by orthogonal trace-less matrices, and calculating the singular values of the matrix $R$.
-It should be mentioned that the singular values of $R$ and hence the value of $\gamma$ do not depend on the bases which we choose for the two parts for expanding the density matrix $\rho$, as long as these bases are orthonormal and traceless.
-Remark: Interestingly, this criterion is a special case of Theorem 1 in [22], where the authors show that this is a powerful criterion and unifies several criteria known before.
-We will now show by several examples that this criterion, originated from the EW defined in (12), is indeed powerful enough to detect entanglement of a class of PPT entangled states which has not been previously detected.
+In this paper, one can use any matrix with any properties to construct entanglement witnesses, as long as eq.
+(12) is satisfied.
+For any given density matrix $\rho$ characterized by the pair $({\bf b},R)$ as shown above, we have to find a witness characterized by a pair $({\bf t},\Lambda)$ such that $\operatorname{Tr}(W\rho)$ as written above is minimized.
+Note that for this kind of witness the parameter ${\bf a}$ in the density matrix plays no role.
+This minimization should be done in the range of all possible entanglement witnesses, namely among all those affine transormations which satisfy the inequality (65), which we rewrite here for convenience:
+$\|{\bf t}\|+\sqrt{d_{1}(d_{1}-1)}\lambda_{max}\leq\sqrt{\frac{d_{2}}{d_{2}-1}}.$ (27)
+To proceed, we need the following:
+Definiiton: A unital entanglement witness is one which corresponds to the Choi matrix of a unital positive map.
+We now show that show that for any given state $\rho$, if a non-unital EW can detect its entanglement, there is a unital EW whcih can detect it in a more effective way.
+This provides a rigorous justification for the focus of researchers in the past years on trying to construct unital positive maps and their corresponding entanglement witnesses.
+Theorem: Given a general bi-partites state $\rho$, unital entanglement witnesses (e.g. those derived from unital maps) are more effective in detecting entanglement of this state.
+Proof: In $\operatorname{Tr}(W\rho)$ given in (26) the vector ${\bf b}$ and the matrix $R$ are determined by the density matrix $\rho$ and our task is to choose the vector ${\bf t}$ and the matrix $\Lambda$ to make $\operatorname{Tr}(W\rho)$ as negative as possible.
+To do so, we invoke the Chauchy Schwartz inequality and choose ${\bf t}$ anti-parallel to ${\bf b}$, that is we take ${\bf t}=-\alpha{\bf b}$, where $\alpha$ is positive constant which can be taken as large as possible provided that the condition (12), namely
+$\|{\bf t}\|+\sqrt{d_{1}(d_{1}-1)}\lambda_{max}\leq\sqrt{\frac{d_{2}}{d_{2}-1}}$
+is not violated.
+To minimize $\operatorname{Tr}(\Lambda R)$, let us use the singular value decomposition of $R$, i.e.
+$R=USV$, where $U,V$ are rotation matrices respectively belonging to orthogonal groups $O(d_{1})$ and $O(d_{2})$ and $S\in M_{d_{1}\times d_{2}}$ is a rectangular matrix on whose main diagonal are the singular values of $R$.
+Then using the freedom of choosing $\Lambda$, we write it as $\Lambda=V^{T}LU^{T}$, where we should choose $L$ such that $\operatorname{Tr}(\Lambda R)$ is minimized subject to the condition (12).
+Thus we have
+$\operatorname{Tr}(\Lambda R)=\operatorname{Tr}(LS).$ (28)
+Noting that the singular values of $R$ (i.e.
+the diagonal entries of $S$) are all positive, this equation can be minimized by considering $L$ to have on its diagonal, a non-negative parameter denoted by $-\beta$ and all the other off-diagonal elements to be zero.
+As explained in detail in appendix I, this leads to $\operatorname{Tr}(R\Lambda)=-\beta\|R\|_{1}.$ We then find
+$\operatorname{Tr}(W\rho)=1-\alpha\|{\bf b}\|^{2}-d_{1}\beta\|R\|_{1}.$ (29)
+We have to find the minimum of this quantity in the region $\alpha-\beta$ defined by
+$\alpha\|{\bf b}\|+\sqrt{d_{1}(d_{1}-1)}\beta\leq\sqrt{\frac{d_{2}}{d_{2}-1}}.$ (30)
+This is now a simple mathematical problem which can be expressed most simply when we cast it as follows: We want to find the minimum value of the function $F(\alpha,\beta):=\operatorname{Tr}(\rho W)$
+$F(\alpha,\beta):=1-p\alpha-q\beta,$
+in the region
+$p^{\prime}\alpha+q^{\prime}\beta\leq 1,$
+where
+$p=\|{\bf b}\|^{2},\ \ \ q=d_{1}\|R\|_{1},\ \ \ p^{\prime}={\|{\bf b}\|}\sqrt{\frac{d_{2}-1}{d_{2}}},\ \ \ q^{\prime}=\sqrt{\frac{d_{1}(d_{1}-1)(d_{2}-1)}{d_{2}}}.$ (31)
+The function $F$ being a linear function of $\alpha$ and $\beta$ aquires its minimum on one of the three vertices of the triangle, figure (2).
+Figure 2: We should minimize the amount of $\operatorname{Tr}(W\rho)=1-p\alpha-q\beta$ with respect to the gray area, where $p,q,p^{\prime},q^{\prime}$ are difined in eq. (31).
+On the vertex $(0,0)$, we have $F(0,0)=1,$ which is positive, hence no detection of entanglement.
+To explore the value of $F$ on the other two vertices, we calculate the value of $F$ on the oblique edge of the triangle defined by $\alpha p^{\prime}+\beta q^{\prime}=1$.
+On this edge, the function $F$ depends on one parameter $\alpha$, which after eliminating $\beta$ in favor of $\alpha$, allows us to define the following one-parameter function $\tilde{F}(\alpha)=\operatorname{Tr}(\rho W)$
+$\tilde{F}(\alpha)=1-p\alpha-q(\frac{1-\alpha p^{\prime}}{q^{\prime}})=1-\frac{q}{q^{\prime}}-\alpha p^{\prime}(\frac{p}{p^{\prime}}-\frac{q}{q^{\prime}}).$ (32)
+Note that the parameters $p,p^{\prime},q$ and $q^{\prime}$ come from the density matrix $\rho$ and $\alpha$ comes from the witness $W$ which is to be chosen so that the value of $\tilde{F}$ is as negative as possible.
+In case that $\frac{p}{p^{\prime}}<\frac{q}{q^{\prime}}$, we take $\alpha$ to be zero leading to the minimum value
+$\tilde{F}_{min}=\tilde{F}(\alpha=0)=1-\frac{q}{q^{\prime}}.$ (33)
+In case that $\frac{p}{p^{\prime}}>\frac{q}{q^{\prime}}$, we take $\alpha$ to be as large as possible, namely $\frac{1}{p^{\prime}}$ which gives the minimum value
+$\tilde{F}_{min}=\tilde{F}(\alpha=\frac{1}{p^{\prime}})=1-\frac{p}{p^{\prime}}.$ (34)
+But the condition $\frac{p}{p^{\prime}}<\frac{q}{q^{\prime}}$ already shows that $\tilde{F}(\alpha=0)\leq\tilde{F}(\alpha=\frac{1}{p^{\prime}})$ which means that there is no advantage in using non-unital witnesses, that is, any entangled state which can be detected by a non-unital witness ($\alpha=\frac{1}{q^{\prime}}$), is also detected by a unital witness ($\alpha=0$).
+Those states which are detected by this unital witness are those for which $\tilde{F}(0)=1-\frac{q^{\prime}}{q}<0$.
+Inserting the values of $q^{\prime}$ and $q$ from (31), this leads to the following condition:
+${\rm If\ \ }\|R\|_{1}>\sqrt{\frac{(d_{1}-1)(d_{2}-1)}{d_{1}d_{2}}},\ \ \ \ \ \ {\rm the\ state\ is\ entangled.}$ (35)
+We stress the significance of this result which is now independent of any entanglemet witnesses and is a criterion based entirely on the density matrix itself.
+We will now show by several examples that this criterion, originated from the EW defined in (24), is indeed powerful enough to detect entanglement of a class of PPT entangled states and indeed in some examples, the range of parameters where the state is shown to be entangled is much larger than previously reported.
 5 Examples
 In this section, we consider a few examples and show the power of our criterion.
-In all the following examples, except example 4.2, the states are PPT entangled states reported in previous works.
+This criterion which is derived here by starting from Mehta’s lemma and positive maps, is a special case of a more general criteria already reported in [21].
+In all the following examples, except example 5.1, the states are PPT entangled states reported in previous works [22, 23].
 We will show that the criterion given in (35), can detect entanglement in a larger interval of the parameters previously reported, and it also can detect PPT entangled states with different dimensions, i.e.
@@
 5.1 Example 1:
-In this example, we demonstrate that the method explained in this paper can identify PPT entangled states, and thus, it can generate indecomposable EWs.
-It has been shown in [23] that the following state is PPT entangled state when $x\neq 1$ (for $x=1$ this is a separable state):
-$\rho=\frac{1}{N}\left(\begin{array}[]{ccc|ccc|ccc}1&.&.&.&1&.&.&.&1\\ .&\frac{1}{x}&.&.&.&.&.&.&.\\ .&.&x&.&.&.&.&.&.\\ \hline\cr.&.&.&x&.&.&.&.&.\\ 1&.&.&.&1&.&.&.&1\\ .&.&.&.&.&\frac{1}{x}&.&.&.\\ \hline\cr.&.&.&.&.&.&\frac{1}{x}&.&.\\ .&.&.&.&.&.&.&x&.\\ 1&.&.&.&1&.&.&.&1\\ \end{array}\right).$ (23)
-Here dots represent zeros, $N=3+x+x^{-1}$ is the normalization factor and $x>0$. The above state can be written in the basis of normalized Gell-Mann matrices as follows:
-[…]
-$+$ $\left(1-\frac{\xi}{2}\right)(\Gamma_{7}\otimes\Gamma_{7}+\Gamma_{8}\otimes\Gamma_{8})-\frac{\sqrt{3}}{2}\eta(\Gamma_{7}\otimes\Gamma_{8}-\Gamma_{8}\otimes\Gamma_{7})\Big{]},$
-where $\xi=x+x^{-1}$ and $\eta=x-x^{-1}$. The eigenvalues of the matrix $R$ are easily found to be
-$\lambda(R)=\frac{1}{N}\{1(3),-1(3),(1-\frac{\xi}{2})\pm i\frac{\sqrt{3}}{2}\eta\},$
-where the numbers in parentheses indicate the multiplicity of each singular value. Thus, taking into account the normalization factor $N$, we find
-$Tr(S)=\frac{6+2\sqrt{(1-\frac{\xi}{2})^{2}+\frac{3}{4}\eta^{2}}}{3(1+\xi)}.$
-Condition (20) for entanglement of the state now amounts to
-$\frac{1}{4}(\xi^{2}-\eta^{2})<\xi-1,$
-or after simplification (inserting $\xi=x+x^{-1}$ and $\eta=x-x^{{}_{1}}$) leads to the following condition for the state (23) to be entangled:
-$2<x+x^{-1}.$
-This shows that the state $\rho_{x}$ is entangled for all values of $x$, except for $x=1$, which coincides with the result of [23].
-4.2 Example 2:
 This example demonstrates that our method can identify the entanglement of all NPT states in the following form in any dimension $d$:
 $\rho=p|\psi\rangle\langle\psi|+(1-p)\frac{I}{d},\quad|\psi\rangle=\frac{1}{\sqrt{d}}\sum_{i=0}^{d-1}|i,i\rangle.$
-To calculate $\gamma$ in equation (20), we need to obtain the coefficients $r_{ij}$:
+It is readily found that
 $r_{ij}=p\operatorname{Tr}[\frac{1}{d}(\sum_{k,l}|kk\rangle\langle ll|).(\Gamma_{i}\otimes\Gamma_{j})]=\frac{p}{d}\operatorname{Tr}(\Gamma_{i}\Gamma_{j}^{T})=\pm\frac{p}{d}\delta_{ij}.$
@@
 Therefore, we find for this state
-$\gamma=\operatorname{Tr}(S)=\frac{p}{d}(d^{2}-1),$
-which, according to eq (20), should be greater than $\frac{d-1}{d}$ so the state is entangled.
+$\|R\|_{1}=\operatorname{Tr}(\sqrt{R^{T}R})=\frac{p}{d}(d^{2}-1),$
+which, according to eq (35), should be greater than $\frac{d-1}{d}$ for the state to be entangled.
 Thus the range of $p$ for the state to be entangled is $p>\frac{1}{d+1}$.
@@
 This result has already been reported in [24].
-4.3 Example 3
-In this example, we provide a PPT entangled state which can be detected by the criterion in equation (20).
-To construct a PPT entangled state we define the following Unextendible Product Basis in $C_{3}\otimes C_{4}:$
-$|\phi_{1}\rangle=\frac{1}{\sqrt{2}}|0\rangle|1-2\rangle,\quad|\phi_{2}\rangle=\frac{1}{\sqrt{6}}|0\rangle(2|0\rangle-|1+2\rangle),\quad|\phi_{3}\rangle=\frac{1}{\sqrt{2}}|1\rangle|1-2\rangle,\quad$
-$|\phi_{4}\rangle=\frac{1}{\sqrt{2}}|2\rangle|1-2\rangle,\quad|\phi_{5}\rangle=\frac{1}{\sqrt{6}}|2\rangle(|1+2\rangle-2|3\rangle),\quad|\phi_{6}\rangle=\frac{1}{\sqrt{2}}|1-2\rangle|0\rangle,\quad$ (25)
-$|\phi_{7}\rangle=\frac{1}{\sqrt{2}}|0-1\rangle|3\rangle,\quad|\phi_{8}\rangle=\frac{1}{\sqrt{12}}|0+1+2\rangle|0+1+2+3\rangle,$ (26)
-and construct the PPT entangled state $\rho=I-\sum_{i}|\phi_{i}\rangle\langle\phi_{i}|$ [8], whose explicit form is given by
-$\rho=\frac{1}{48}\left(\begin{array}[]{c|c|c|c}3\ \epsilon&-\epsilon&-\epsilon&-\epsilon\\ \hline\cr-\epsilon&D&D^{\prime}&-\epsilon\\ \hline\cr-\epsilon&D^{\prime T}&D&-\epsilon\\ \hline\cr-\epsilon&-\epsilon&-\epsilon&3\ \epsilon\\ \end{array}\right),$
-where $\epsilon$ is the 3 dimensional square matrix all of its elements are equal to 1, and
-$D=\left(\begin{array}[]{ccc}5&-1&-1\\ -1&5&-1\\ -1&-1&5\end{array}\right),\quad D^{\prime}=\left(\begin{array}[]{ccc}-1&5&-1\\ -1&-1&5\\ 5&-1&-1\end{array}\right).$
-The matrix $R$ is simply obtained by noting that its elements are given by $r_{mn}=\operatorname{Tr}(\rho(\Gamma_{m}\otimes\Omega_{n}))=-\sum_{i}\langle\phi_{i}|\Gamma_{m}\otimes\Omega_{n}|\phi_{i}\rangle$.
-The result is
-$R=\frac{1}{24}\left(\begin{array}[]{c|c|c}-\epsilon_{3\times 6}&0_{3\times 6}&A_{3\times 3}\\ \hline\cr 0_{3\times 6}&0_{3\times 6}&0_{3\times 3}\\ \hline\cr B_{2\times 6}&0_{2\times 6}&C_{2\times 3}\end{array}\right),$
-where $\epsilon_{3\times 6}$ is a $3\times 6$ matrix all of whose elements are 1, and similarly ${0}_{3\times 6}$ is a matrix all of whose elements are zero. Also
-$A$ $=$ $\left(\begin{array}[]{ccc}0&0&-3\sqrt{\frac{3}{2}}\\ 0&0&0\\ 3&\sqrt{3}&\sqrt{\frac{3}{2}}\end{array}\right),$ (30)
-$B$ $=$ $\left(\begin{array}[]{cccccc}{2}&{2}&-1&0&0&0\\ \frac{2}{\sqrt{3}}&\frac{2}{\sqrt{3}}&\frac{1}{\sqrt{3}}&0&\frac{-4}{\sqrt{3}}&\frac{-4}{\sqrt{3}}\end{array}\right),$ (33)
-$C$ $=$ $\left(\begin{array}[]{ccc}0&0&-\frac{1}{2}\sqrt{\frac{3}{2}}\\ -\frac{1}{\sqrt{3}}&-\frac{1}{3}&-\frac{5}{6\sqrt{2}}\end{array}\right).$ (36)
-The singular values of this matrix are obtained numerically and one finds that
-$\gamma=\operatorname{Tr}(S)=1.429>\sqrt{\frac{3\times 4}{2\times 3}}=\sqrt{2},$
-which shows that while $\rho$ is a PPT entangled state, it can be detected by the criterion (20).
 5.2 Example 2
@@
 This leads to
-$\gamma=\frac{1}{4}\frac{(a-1)^{2}}{(a+1)^{2}}+3\frac{|a|}{(a+1)^{2}}+\frac{1}{2}\left|\frac{a-1}{a+1}\right|.$
+$\|R\|_{1}=\operatorname{Tr}(\sqrt{R^{T}R})=\frac{1}{4}\frac{(a-1)^{2}}{(a+1)^{2}}+3\frac{|a|}{(a+1)^{2}}+\frac{1}{2}\left|\frac{a-1}{a+1}\right|.$
 Figure 3 shows that the value of $\gamma$ is always larger than $\sqrt{\frac{(d_{1}-1)(d_{2}-1)}{d_{1}d_{2}}}=\frac{3}{4}$ except for $a=1$.
@@
 which is obviously separable. Incidentally, this is the limiting case where $\gamma$ approaches the value $\frac{3}{4}$.
-4.5 Example 5
-This example is a continuation of the previous example and demonstrates that for the second PPT entangled state introduced in [23], the interval of entanglement (indicated as $10<b<$ 1), is indeed much larger and as in the previous example covers all positive and finite values of $b$ different from $1$.
-The state is given by
-$\rho=\frac{1}{N}\left(\begin{array}[]{cccc|cccc|cccc|cccc}1&.&.&.&.&1&.&.&.&.&1&.&.&.&.&1\\ .&\frac{1}{b}&.&.&1&.&.&.&.&.&.&.&.&.&.&.\\ .&.&1&.&.&.&.&.&1&.&.&.&.&.&.&.\\ .&.&.&b&.&.&.&.&.&.&.&.&1&.&.&.\\ \hline\cr.&1&.&.&b&.&.&.&.&.&.&.&.&.&.&.\\ 1&.&.&.&.&1&.&.&.&.&1&.&.&.&.&1\\ .&.&.&.&.&.&\frac{1}{b}&.&.&1&.&.&.&.&.&.\\ .&.&.&.&.&.&.&1&.&.&.&.&.&1&.&.\\ \hline\cr.&.&1&.&.&.&.&.&1&.&.&.&.&.&.&.\\ .&.&.&.&.&.&1&.&.&b&.&.&.&.&.&.\\ 1&.&.&.&.&1&.&.&.&.&1&.&.&.&.&1\\ .&.&.&.&.&.&.&.&.&.&.&\frac{1}{b}&.&.&1&.\\ \hline\cr.&.&.&1&.&.&.&.&.&.&.&.&\frac{1}{b}&.&.&.\\ .&.&.&.&.&.&.&1&.&.&.&.&.&1&.&.\\ .&.&.&.&.&.&.&.&.&.&.&1&.&.&b&.\\ 1&.&.&.&.&1&.&.&.&.&1&.&.&.&.&1\\ \end{array}\right),$
-where $N=4b+\frac{4}{b}+8$, and its corresponding matrix is
-$R=\frac{1}{(1+b)^{2}}\left(\frac{b}{2}I_{6}\oplus 0_{6}\oplus A\right),$
-where $I_{6}$ and $0_{6}$ are respectively the identity matrix and the zero matrix in $C_{6}$ and
-$A=\left(\begin{array}[]{ccc}-\frac{1}{8}(b-1)^{2}&-\frac{b^{2}+2b-3}{8\sqrt{3}}&-\frac{(b-1)b}{2\sqrt{6}}\\ \frac{3b^{2}-2b-1}{8\sqrt{3}}&-\frac{1}{24}(b-1)^{2}&-\frac{b^{2}+b-2}{6\sqrt{2}}\\ \frac{b-1}{2\sqrt{6}}&\frac{2b^{2}-b-1}{6\sqrt{2}}&-\frac{1}{12}(b-1)^{2}\\ \end{array}\right),$
-with singular values given as
-$\Big{\{}\frac{\left|b-1\right|}{4\left|b+1\right|}\ (2),\ \frac{\left|b\right|}{2(b+1)^{2}}\ (6),\ \frac{(b-1)^{2}}{4(b+1)^{2}}\ (1)\Big{\}},$
-which is the same as the previous example and gives the same result.
 5.3 Example 3
@@
 and we find
-$\gamma=\frac{1}{2\sqrt{10}}+\frac{4}{5\sqrt{2}}=\big{(}\frac{8+\sqrt{5}}{10}\big{)}\frac{1}{\sqrt{2}},$
+$\|R\|_{1}=\operatorname{Tr}(\sqrt{R^{T}R})=\frac{1}{2\sqrt{10}}+\frac{4}{5\sqrt{2}}=\big{(}\frac{8+\sqrt{5}}{10}\big{)}\frac{1}{\sqrt{2}},$
 which is greater than $\sqrt{\frac{(d_{1}-1)(d_{2}-1)}{d_{1}d_{2}}}=\frac{1}{\sqrt{2}}$. This shows that the PPT state (57) is entangled.
 6 Conclusion
-In this paper, a class of Entanglement Witnesses for two particle systems with different dimensions was presented.
-Specifically we showed in section (3), how the use of this class of EW’s can be reduced to a simple criterion, (equation (20)), which can be checked on the form of the density matrix.
+Mehta’s lemma is a simple sufficient condition for a matrix to be positive.
+We have used this lemma to construct a large family of positive maps and characterized them in terms of their corresponding affine transformations, i.e.
+transformation of the so-called Bloch vector ${\bf r}\longrightarrow\Lambda{\bf r}+{\bf t}$, equation (12).
+Using these positive maps, and by using the one-to-one correspondence between entanglement witnesses and the Choi matrix of positive maps, a multi-parameter family of Entanglement Witnesses for two particle systems with different dimensions was constructed.
+Specifically we showed how the use of this class of EW’s can be reduced to a simple criterion, (equation (35)), which can be checked on the form of the density matrix.
 We then showed by checking several one-parameter family of PPT entangled states that this criterion can detect their entanglement in a region of parameters much larger than previously thought.
+We also rigorously proved that unital entanglement witnesses, i.e.
+those corresponding to unital positive maps, can detect a larger set of entangled states.
+Acknowledgement: This research was supported in part by Iran National Science Foundation, under Grant No.4022322.
+The authors wish to thank S.
+Roofeh, A.
+Farmanian, and A.
+Tangestani, for many inspiring discussions during this project.
+The authors also thank Dariusz Chruscinski for his valuable comments.